### v1

# DistPred: A Distribution-Free Probabilistic Inference Method for Regression and Forecasting

Abstract
Traditional regression and prediction tasks often only provide deterministic point estimates. To estimate the uncertainty or distribution information of the response variable, methods such as Bayesian inference, model ensembling, or MC Dropout are typically used. These methods either assume that the posterior distribution of samples follows a Gaussian process or require thousands of forward passes for sample generation. We propose a novel approach called DistPred for regression and forecasting tasks, which overcomes the limitations of existing methods while remaining simple and powerful. Specifically, we transform proper scoring rules that measure the discrepancy between the predicted distribution and the target distribution into a differentiable discrete form and use it as a loss function to train the model end-to-end. This allows the model to sample numerous samples in a single forward pass to estimate the potential distribution of the response variable. We have compared our method with several existing approaches on multiple datasets and achieved state-of-the-art performance. Additionally, our method significantly improves computational efficiency. For example, compared to state-of-the-art models, DistPred has a 90x faster inference speed. Experimental results can be reproduced through this Repository.
1 Introduction
Traditional deterministic point estimates are no longer sufficient to meet the needs of AI safety and uncertainty quantification. For example, we may want to obtain confidence intervals for predicted points to make important decisions, such as deciding whether to travel based on weather forecasts or how to invest based on stock predictions. Moreover, this is particularly important in high-security AI application areas such as autonomous driving, risk estimation, and decision-making.
In this paper, we consider the underlying distribution behind predicting the response variable because it reflects the confidence intervals of all levels. For example, based on this distribution, we can calculate confidence intervals at any level, coverage rate, and uncertainty quantification. Currently, predicting the distribution of the response variable poses a challenge because at a specific moment, the response variable can only take on a single deterministic value. This point can be viewed as a maximum likelihood sample from its underlying distribution, but it fails to reflect the overall state of the underlying distribution.
Currently, the primary method used to address distribution prediction and uncertainty quantification in regression and prediction tasks is frequentist sampling. These methods involve sampling numerous samples by perturbing the explanatory variable or model to approximate the underlying distribution of the response variable. For instance, Bayesian neural networks (BNNs) simulate this uncertainty by assuming that their parameters follow a Gaussian distribution, thereby capturing the model’s uncertainty given the data (Blundell et al., 2015). Similarly, ensemble-based methods have been proposed to combine multiple deep models with random outputs to capture prediction uncertainty. MC Dropout (Gal and Ghahramani, 2016) demonstrates that enabling dropout during each testing process yields results akin to model ensembling. Additionally, models based on GANs and diffusion are introduced for conditional density estimation and prediction uncertainty quantification. These models utilize noise during the generation or diffusion process to obtain different predicted values for estimating the uncertainty of the response variable.
The common characteristic of these methods mentioned above is the requirement of $K$ forward passes to sample $K$ representative samples. For example, Bayesian framework-based methods require $K$ learnable parameter samples to be inferred to obtain $K$ representative samples; ensemble methods require $K$ models to jointly infer; MC Dropout requires $K$ forward passes with random dropout activation; generative models require $K$ forward or diffusion processes. However, the excessive forward passes result in significant computational overhead and slow speed, a drawback that becomes increasingly apparent for AI applications with high real-time requirements.
Furthermore, several methods (Nix and Weigend, 1994; Salinas et al., 2020; He et al., 2020) transform distribution prediction and uncertainty quantification into predicting statistical variables such as mean and variance by assuming that the response variable follows a known continuous distribution. For instance, DeepAR (Salinas et al., 2020) assumes the response variable follows a Gaussian distribution, thereby directly utilizing GaussianNLLLoss (Nix and Weigend, 1994) to optimize its mean and variance. These methods only predict statistical variables, reducing the inference cost, but strong distribution assumptions often fail to capture the true data distribution, resulting in inferior performance.
Additionally, there are methods that train models using quantiles as loss functions to obtain specific quantile intervals (QIs) of the response variable. For example, to predict the 5th and 95th percentile confidence intervals, one can directly predict the corresponding quantiles. MRQNN predicts specified quantiles using quantile loss. Similarly, TFT enables the model to directly output three quantiles: 0.1, 0.5, and 0.9. These methods cannot provide distribution information about the response variable but only offer fixed QIs, thus lacking flexibility in their predictions.
Figure 1: DistPred can provide $K$ predicted values $\hat{y}$ of the response variable $y$ given the predictor variable $x$ in a single forward process, denoted as $\mathbb{E}(\hat{Y}|x)$, where $\hat{Y}$ represents a maximum likelihood sample of $y$. Based on this sampling, the probability mass distribution (PMD) $P_{\theta}(\hat{y}|x)$, cumulative distribution function (PDF) $F_{\theta}(\hat{y}|x)$, and confidence curve $CC_{\theta}(\hat{y}|x)$ for the response variable $y$ can be computed, thereby yielding comprehensive statistical insights into $y$. For instance, this includes confidence intervals (CI) at any desired level, as well as p-values.
To address this issue, we propose a novel method called DistPred, which is a distribution-free probabilistic inference method for regression and forecasting tasks. DistPred is a simple and powerful method that can estimate the distribution of the response variable in a single forward pass. Specifically, we contemplate employing all predictive quantiles to specify the potential cumulative density function (CDF) of the predictor variable, and we show that the full quantiles’ prediction can be translated into calculating the minimum expected score of the response variable and the predictive ensemble variables. Based on this, we transform proper scoring rules that measure the discrepancy between the predicted distribution and the target distribution into a differentiable discrete form and use it as a loss function to train the model end-to-end. This allows the model to sample numerous samples in a single forward pass to estimate the potential distribution of the response variable. DistPred is orthogonal to other methods, enabling its combination with alternative approaches to enhance estimation performance. Further, we show that DistPred can provide comprehensive statistical insights into the response variable, including confidence intervals at any desired level, p-values, and other statistical information, as shown in Figure 1. Experimental results show that DistPred outperforms existing methods in terms of both accuracy and computational efficiency. Specifically, DistPred has a 90x faster inference speed than state-of-the-art models.
2 Method
Assume that the dataset $D=\{x_{i},y_{i}\}_{i=1}^{N}$ consists of $N$ sample-label pairs. The subscript $i$ will be omitted if it does not cause ambiguity in the context. Our objective is to utilize a machine learning model $M$ with parameters $\theta$ to predict the underlying distribution $P(y)$ of the response variable $y$ from $D$, aiming to acquire comprehensive statistical insights such as obtaining confidence intervals (CI) and quantifying uncertainty at any desired level.
Direct prediction of distribution $P$ is not feasible, because:
  1. 1.
    Without distributional assumptions, we cannot give a valid representation of the PDF or CDF of the predicted distribution $P_{\theta}(\hat{y})$.
  2. 2.
    We can only obtain a singular deterministic value for the response variable $y$, lacking access to its distributional information to guide model learning.
Figure 2: Relationship between CDF and all predictive quantiles.
To address the aforementioned issues, we contemplate employing all predictive quantiles $\hat{q}_{1},\hat{q}_{2},\cdots$, at levels $\alpha_{1},\alpha_{2},\cdots$ to specify the potential CDF $F_{\theta}(\hat{y})$ of the predictor variable $\hat{y}$. This is because if we know the cumulative distribution function of a random variable, we can find any quantile by setting $F(y)=q$. Conversely, if we have a complete set of quantiles, we can approximate or reconstruct the cumulative distribution function of the random variable. As shwon in Figure 2, quantiles provide discrete ‘snapshots’ of the distribution, while the CDF is a continuous, smooth version of these snapshots, offering a complete description of the cumulative probability from the minimum to the maximum value.
Next, we will present that the full predictive quantiles outlined earlier serves as a proper approximation of the response variable’s CDF. And the full quantiles’ prediction can be translated into calculating the minimum expected score of the response variable and the predictive ensemble variables $\hat{Y}$, where $\hat{Y}=\{\hat{y}_{1},\cdots,\hat{y}_{K}\}$. Prior to delving into this analysis, we shall initially introduce the scoring rules utilized for assessing the propriety of predicted distributions.
2.1 Utilize proper scoring rules as the loss function
Scoring rules offer a concise measure for assessing probabilistic forecasts by assigning numerical scores according to the forecast distribution and predicted outcomes (Gneiting and Raftery, 2007; Jordan et al., 2017). Specifically, Let $\Omega$ denote the set of possible values of the quantity of interest, and let $\mathcal{P}$ denote a convex class of probability distributions on $\Omega$. A scoring rule is a function
$S:\Omega\times\mathcal{P}\rightarrow\mathbb{R}\cup\{\infty\}$ (1)
that assigns numerical values to pairs of forecasts $P\in\mathcal{P}$ and observations $y\in\Omega$. We identify probabilistic forecasts $P$ with the associated CDF $F$ or PDF $f$, and consider scoring rules to be negatively oriented, where a lower score signifies a more accurate forecast. A proper scoring rule is optimized when the forecast aligns with the true distribution of the observation, i.e., if
$E_{Y\sim Q}[S(Q,Y)]\leq E_{Y\sim Q}[S(P,Y)]$ (2)
for all $P,Q\in\mathcal{P}$. A scoring rule is termed strictly proper when equality is achieved only when $P=Q$. Proper scoring rules (PSR) are essential for comparative evaluation, particularly in ranking forecasts. In practice, the forecaster with the lowest average score across multiple forecast cases typically demonstrates the best predictive performance. Proper scoring rules incentivize forecasters to accurately report their perception of the true distribution in this scenario. Therefore, PSR provide attractive loss and utility functions that can be tailored to a regression or forecast problem. To estimate $\theta$, we might measure the goodness-of-fit by the mean score
$S_{n}(\theta)=\frac{1}{N}\sum_{i=1}^{N}S(P_{\theta}(\hat{y}_{i}),y_{i}).$ (3)
Let $\theta^{*}$ denotes the true parameter value, then asymptotic arguments indicate that $\text{argmin}_{\theta}S_{n}(\theta)\rightarrow\theta^{*}$ as $n\rightarrow\infty$. This suggests a general approach to transform PSR into loss functions for training models, which implicitly minimizes the divergence between predictive and true distributions.
2.2 Property of full predictive quantiles
We contemplate probabilistic predictions pertaining to a continuous quantity, manifested as full predictive quantiles $\hat{q}_{1},\cdots,\hat{q}_{K}$. For $P\in\mathcal{P}$, let $q_{1},\cdots,q_{K}$ denote the true $P$-quantiles at levels $\alpha_{1},\cdots,\alpha_{K}\in(0,1)$. Then, the expected score $S(q_{1},\cdots,q_{K};P)$ can be defined as
$S(\hat{q}_{1},\cdots,\hat{q}_{K};P)=\int S(\hat{q}_{1},\cdots,\hat{q}_{K};y)\,\text{d}P(y).$ (4)
Further, a scoring rule S is proper if $S(q_{1},\cdots,q_{K};P)\geq S(\hat{q}_{1},\cdots,\hat{q}_{K};P)$. Based on this definition, we assuming that $s_{k},i\in[1,\cdots K]$ is nondecreasing and $h$ is arbitrary, then the scoring rule
$S(\hat{q}_{1},\cdots,\hat{q}_{K};P)=\sum_{k=1}^{K}\left(\alpha_{i}s_{k}(\hat{q}_{k})+(s_{k}(y)-s_{k}(\hat{q_{k}})\mathbbm{1}\{y\leq\hat{q}_{k}\})\right)$ (5)
is proper for predicting the quantiles at levels $\alpha_{1},\cdots,\alpha_{K}$ when $K\rightarrow\infty$ (Gneiting and Raftery, 2007; Schervish et al., 2012). $\mathbbm{1}\{y\leq\hat{q}_{k}\}$ denotes the indicator function which is one if $y\leq\hat{q}_{k}$ and zero otherwise.
Equation 5 shows that full predictive quantiles is proper. In essence, delineating a predictive CDF is equivalent to specifying all predictive quantiles. Consequently, we can formulate scoring rules for the predictive distribution based on the scoring rules for the quantiles. Specifically, let $S_{\alpha}$ denotes a proper scoring rule for the quantile at level $\alpha$, then the scoring rule
$S(F,y)=\int_{0}^{1}S_{\alpha}(F^{-1}(\alpha);y)\,\text{d}\alpha=\int_{-\infty}^{\infty}S(F(\hat{y}),\mathbbm{1}\{y\leq\hat{y}\})\,\text{d}\hat{y}$ (6)
is proper. Here, we can find that the right of Equation 6 corresponds to the CRPS in which $S$ is the quadratic or Brier score, which is defined as
$C(F,y)=\int_{-\infty}^{\infty}(F(\hat{y})-\mathbbm{1}\{y\leq\hat{y}\})^{2}\,\text{d}\hat{y}.$ (7)
If the first moment of $F$ is finite, the CRPS can be written as
$C(F,y)=\mathbb{E}_{F}[|\hat{Y}-y|]-\frac{1}{2}\mathbb{E}_{FF}[|\hat{Y}-\hat{Y}^{\prime}|],$ (8)
where $\hat{Y}$ and $\hat{Y}^{\prime}$ denote independent predictive variables with distribution $F$.
Figure 3: The workflow of DistPred. An ensemble of predictive variables $\hat{Y}$ is inferred in a forward pass and PSR $S(\mathbb{E}(\hat{Y}|x),y)$ is utilized to train the learner end-to-end.
2.3 End-to-end ensemble inference
Figure 4: DistPred’s architecture.
Based on the analysis provided above, it is evident that predicting the full quantiles is equivalent to minimizing Equation 8 w.r.t $\mathbb{E}(\hat{Y}|y)$. Hence, as the workflow shown in Figure 3, we can develop a model $M$ with parameters $\theta$ that infers an ensemble of predictive variables $\hat{Y}$ in a forward pass and utilize Equation 8 to train it end-to-end. This allows the model to sample numerous samples in a single forward pass to estimate the empirical CDF $\hat{F}$ by the predictive ensemble variables
$C(\hat{F},y)=\frac{1}{K}\sum_{k=1}^{K}|\hat{y}_{k}-y|-\frac{1}{2K^{2}}\sum_{k=1}^{K}\sum_{j=1}^{K}|\hat{y}_{k}-\hat{y}^{\prime}_{j}|.$ (9)
It’s worth noting that Equation 9 is a differentiable discrete form w.r.t $\hat{Y}$ and $\hat{Y^{\prime}}$ that strictly satisfies PSR. However, implementations of Equation 9 exhibit inefficiency due to their computational complexity of $O(K^{2})$. This can be enhanced by employing representations based on the generalized quantile function (Laio and Tamea, 2007) and the sorted predictive ensemble variables
$C(\hat{F},y)=\frac{1}{2K^{2}}\sum_{k=1}^{K}(\overrightarrow{y_{k}}-y)(k\mathbbm{1}\{y\leq\overrightarrow{y_{k}}\}-i+\frac{1}{2}).$ (10)
Since a sorting operation is involved, the computational complexity of Equation 10 is $O(KlogK)$. To conserve memory, we also suggest utilizing Equation 10 as the loss function, as predicting ensemble variables in long-term forecasting tasks may lead to out-of-memory issues on GPUs.
2.4 Incorporate alternative methodologies
DistPred is orthogonal to other methods, enabling its combination with alternative approaches to enhance estimation performance. Here, with a focus on computational efficiency and memory conservation, we opt to integrate MC Dropout with DistPred, thereby denoting the amalgamation as DistPred-MCD. In our experiments, we observed that DistPred-MCD can further enhance uncertainty quantification performance, albeit with a marginal increase in computational effort.
3 Experiments
In this paper, our focus centers on regression (proposed by Hernández-Lobato and Adams (2015)) and prediction (proposed by Zhou et al. (2021a)) tasks, where we validate the application of the proposed DistPred method to these specific endeavors.
3.1 PICP and QICE metrics
In recent years, BNNs (Hernández-Lobato and Adams, 2015; Gal and Ghahramani, 2016) have emerged as a class of models that aims at estimating the uncertainty providing a more complete picture of $p(y|x)$. In (Han et al., 2022), it is contended that the metric utilized in BNNs to assess uncertainty estimates, namely the negative log-likelihood (NLL), is computed based on Gaussian densities. This assumption implies that they consider both the conditional distribution $p(y|x=x^{\prime})$ for all $x^{\prime}$ are Gaussian. However, this assumption is very difficult to verify for real-world datasets. We follow CARD (Han et al., 2022) and use the following two metrics, both of which are designed to empirically evaluate the degree of similarity between learned and true conditional distributions:
  • PICP (Prediction Interval Coverage Probability) (Yao et al., 2019) is a metric that measures the proportion of true labels that fall within the prediction interval.
  • QICE (Quantile Interval Calibration Error) (Han et al., 2022) is a metric that measures the average difference between the predicted and true quantiles at a given level $\alpha$.
The PICP is calculated as
$PICP:=\frac{1}{N}\sum_{n=1}^{N}\mathbbm{1}\{\hat{y}_{n}\geq q_{\alpha/2}\}\cdot\mathbbm{1}\{\hat{y}_{n}\leq q_{1-\alpha/2}\},$ (11)
where $q_{\alpha/2}$ and $q_{1-\alpha/2}$ represent the low and high percentiles, respectively, that we have selected for the predicted $\hat{y}$ outputs given the same $x$ input. This metric evaluates the proportion of accurate observations that lie within the percentile range of the generated $\hat{y}$ samples corresponding to each $x$ input. Within this study, we opt for the $2.5$th and $97.5$th percentiles, signifying that an optimal PICP value for the acquired model would ideally reach 95%.
However, a caveat of the PICP metric becomes apparent in the measurement of distribution differences. For instance, consider a situation where the 2.5th to 97.5th percentile range of the learned distribution encompasses the data falling between the 1st and 96th percentiles of the true distribution. Even with a sufficient number of samples, the resulting PICP value would still approach 95%. However, it becomes evident that a discrepancy exists between the learned distribution and the true distribution.
Drawing from this reasoning, Han et al. (2022) introduces a novel empirical metric called QICE. This metric can be perceived as an enhanced version of PICP, offering finer granularity and addressing the issue of uncovered quantile ranges. To calculate QICE, the initial step involves generating an adequate number of samples for each $\hat{y}$ value. These samples are then divided into $M$ bins of approximately equal sizes. Subsequently, the quantile values are determined at each boundary within these bins. The definition of QICE entails computing the mean absolute error (MAE) between the proportion of true data encompassed by each quantile interval and the optimal proportion, which is 1/M for all intervals:
$QICE:=\frac{1}{M}\sum_{m=1}^{M}|r_{m}-\frac{1}{M}|,\quad\text{where}\ \ r_{m}=\frac{1}{N}\sum_{n=1}^{N}\mathbbm{1}\{\hat{y}_{n}\geq q_{\alpha/2}\}\cdot\mathbbm{1}\{\hat{y}_{n}\leq q_{1-\alpha/2}\}.$ (12)
In this paper, we followed Han et al. (2022) and set $M=10$ for all experiments.
3.2 Toy examples
To demonstrate the effectiveness of DistPred, we initially conduct experiments on $8$ toy examples as done in CARD (Han et al., 2022). These examples are specifically crafted with distinct statistical characteristics in their data generating functions: some have a uni-modal symmetric distribution for their error term (linear regression, quadratic regression, sinusoidal regression), others have heteroscedasticity (log-log linear regression, log-log cubic regression) or multi-modality (inverse sinusoidal regression, 8 Gaussians, full circle).
The research demonstrates that a trained DistPred models has the capability to produce samples that closely resemble the true response variable for novel covariates. Additionally, it can quantitatively match the true distribution based on certain summary statistics. The study visualizes scatter plots comparing real and generated data for all eight tasks in Figure 5. In cases where the tasks involve unimodal conditional distributions, the interest region fills the region between the $2.5$th and $97.5$th percentiles of the generated $\hat{y}$ values.
We note that within every task, the generated samples seamlessly integrate with the authentic test instances, indicating the potential of DistPred to reconstruct the inherent data generation process. This experiment visually demonstrates that DistPred effectively reconstructs the sample potential distribution of the target response variable. This indicates that the advantages of DistPred mentioned earlier can be fully harnessed in distribution prediction.
Figure 5: Scatter plot of DistPred’s regression results on $8$ toy examples.
3.3 UCI regression tasks
For experiments conducted on real-world datasets, we utilize the same $10$ UCI regression benchmark datasets (Asuncion and Newman, 2007) and follow the experimental protocol introduced by Hernández-Lobato and Adams (2015), which has also been followed by Gal and Ghahramani (2016) and Lakshminarayanan et al. (2017), as well as by Han et al. (2022). The dataset information can be found in Table 5 located in Appendix B.
[…]
Dataset PBP MC Dropout Deep Ensembles GCDS CARD DistPred DistPred-MCD
Boston $3.50\pm 0.88$ $3.82\pm 0.82$ $\underline{3.37\pm 0.00}$ $11.73\pm 1.05$ $3.45\pm 0.83$ $\underline{3.22\pm 0.86}$ $\bm{2.95\pm 0.78}$
Concrete $2.52\pm 0.60$ $4.17\pm 1.06$ $2.68\pm 0.64$ $10.49\pm 1.01$ $\underline{2.30\pm 0.66}$ $\underline{2.42\pm 0.43}$ $\bm{2.16\pm 0.43}$
Energy $6.54\pm 0.90$ $5.22\pm 1.02$ $\underline{3.62\pm 0.58}$ $7.41\pm 2.19$ $4.91\pm 0.94$ $\underline{3.73\pm 0.71}$ $\bm{3.39\pm 1.09}$
Kin8nm $1.31\pm 0.25$ $1.50\pm 0.32$ $1.17\pm 0.22$ $7.73\pm 0.80$ $\bm{0.92\pm 0.25}$ $\underline{1.11\pm 0.28}$ $\underline{1.00\pm 0.24}$
Naval $4.06\pm 1.25$ $12.50\pm 1.95$ $6.64\pm 0.60$ $5.76\pm 2.25$ $\bm{0.80\pm 0.21}$ $\underline{3.30\pm 0.91}$ $\underline{1.42\pm 0.47}$
Power $\underline{0.82\pm 0.19}$ $1.32\pm 0.37$ $1.09\pm 0.26$ $1.77\pm 0.33$ $\underline{0.92\pm 0.21}$ $0.97\pm 0.30$ $\bm{0.80\pm 0.23}$
Protein $1.69\pm 0.09$ $2.82\pm 0.41$ $2.17\pm 0.16$ $2.33\pm 0.18$ $\underline{0.71\pm 0.11}$ $\bm{0.47\pm 0.11}$ $\underline{0.52\pm 0.14}$
Wine $\underline{2.22\pm 0.64}$ $2.79\pm 0.56$ $2.37\pm 0.63$ $3.13\pm 0.79$ $3.39\pm 0.69$ $\underline{2.01\pm 0.51}$ $\bm{1.79\pm 0.45}$
Yacht $6.93\pm 1.74$ $10.33\pm 1.34$ $7.22\pm 1.41$ $\underline{5.01\pm 1.02}$ $8.03\pm 1.17$ $\underline{5.31\pm 0.96}$ $\bm{4.37\pm 1.15}$
Year $2.96\pm$ NA $2.43\pm$ NA $2.56\pm$ NA $1.61\pm$ NA $\underline{0.53\pm\text{NA}}$ $\underline{0.58\pm\text{NA}}$ $\bm{0.28\pm\text{NA}}$
# Top 1 00 00 00 00 $2$ $1$ $\bm{7}$
# Top 2 $2$ 00 $2$ $1$ $6$ $9$ $10$
We compare DistPred with other state-of-the-art methods, including PBP (Hernández-Lobato and Adams, 2015), MC Dropout (Gal and Ghahramani, 2016), DeepEnsemble (Lakshminarayanan et al., 2017), and another deep generative model that estimates a conditional distribution sampler, GCDS (Zhou et al., 2021b), as well as diffusion model, CARD (Han et al., 2022). The multiple train-test splits are applied with a 90%/10% ratio, following the same methodology as Hernández-Lobato and Adams (2015) and Han et al. (2022) (20 folds for all datasets except 5 for Protein and 1 for Year). The reported metrics are presented as the mean and standard deviation across all splits. As pointed out by Han et al. (2022), we compare the QICE of different methods on various UCI datasets. Additional information regarding the experimental setup for these models is available in Appendix C. The experimental results, along with corresponding metrics, are presented in Table 1. The frequency with which each model achieves the best corresponding metric is reported in the penultimate row, while the frequency with which it achieves the top two positions is reported in the last row.
The results demonstrate that the DistPred method outperforms existing methods, often by a considerable margin. It is worth noting that these impressive results are achieved in a single forward pass of the DistPred method. Crucially, the performance of uncertain quantization can be enhanced even further by leveraging DistPred-MCD, a hybrid approach combining DistPred and MC Dropout.
The implementation of DistPred on UCI regression tasks follows a straightforward approach: We employ a basic multilayer perceptron (MLP) as the foundational framework, complemented by Equation 9 serving as the loss function for end-to-end training. Due to the fact that DistPred necessitates solely a single forward inference, its inference speed is notably rapid. Table 2 presents a comparison of the training and inference speeds of mainstream models. It should be noted that, for a fair comparison, the implementations of various models are constructed on the same backbone and utilized the same equipment. It is evident that DistPred is approximately $230$ times faster in training and about $90$ times faster in inference compared to the state-of-the-art model CARD. The inference speed of DistPred is slower than its training speed because it involves calculating distribution statistical metrics like QICE and PICP.
Table 2: Comparison of model training and inference times (minutes) on UCI boston datasets.
Models DistPred DistPred-MCD PBP MC Dropout CARD
Training 0.035 $\pm$ 6E-3 0.035 $\pm$ 6E-3 0.04 $\pm$ 8E-3 0.031 $\pm$ 0.01 8.14$\pm$ 0.05
Inference 0.095 $\pm$ 4E-3 0.027 $\pm$ 5E-3 5.23 $\pm$ 0.10 4.62 $\pm$ 0.06 8.31 $\pm$ 0.17
3.4 Ablation study of the number of samples and ensembles
We investigate the influence of the number of samples generated by DistPred, as well as the number of ensembles of DistPred-MCD, on their respective performances. As shown in Figure 6, with an increase in the number of output samples and ensembles, the model’s performance shows a gradual improvement, eventually reaching a point of saturation.
Figure 6: Ablation study of the number of samples in DistPred (a) and the number of ensembles in DistPred-MCD (b).
3.5 Time series distribution forecasting
We extend time series forecasting (Zhou et al., 2021a; Wu et al., 2021; Zhou et al., 2022; Liu et al., 2023) from point estimation to the task of distribution prediction to infer about more statistical information about a certain moment.
Baselines: We employ recent 10 SOTA methods for comparisons, including iTransformer Liu et al. (2023), PatchTST Nie et al. (2022), SCINet Liu et al. (2022a), TimesNet Wu et al. (2022), DLinear Zeng et al. (2023), FEDformer Zhou et al. (2022), Autoformer Wu et al. (2021), Informer Zhou et al. (2021a), LogTrans Li et al. (2019) and Reformer Kitaev et al. (2020). We use the same experimental setup as (Zhou et al., 2021a) and (Liu et al., 2022a) and follow the same experimental protocol as (Zhou et al., 2021a). Univariate results can be found in Appendix D.
Datasets and setting: The detailed information pertaining to the datasets can be located in Appendix B. The models used in the experiments are evaluated over a wide range of prediction lengths to compare performance on different future horizons: 96, 192, 336 and 720. The experimental settings are the same for both multivariate and univariate tasks. We use the average of the MSE and MAE ($\frac{MSE+MAE}{2}$) to evaluate the overall performance of the model. It is noteworthy that DistPred provides an ensemble $\hat{Y}$ of response variables. Consequently, we employ the mean value of $\hat{Y}$ as the point estimate at that moment.
Main results: The results for multivariate TS forecasting are outlined in Table 3, with the optimal results highlighted in bold and the second-best results emphasized with underlined. It can be found that, despite not utilizing MSE and MAE, DistPred achieves state-of-the-art performance across all datasets and prediction length configurations. iTransformer and PatchTST stand out as the latest models acknowledged for their exceptional average performance. Compared with them, the proposed DistPred demonstrates an average performance increase of 3.5% and 16.5%, respectively, achieving a substantial performance improvement. We provide metrics, e.g., CRPS, QICE, PICP, for comparison by the future research community.
Table 3: Multivariate time series forecasting results on six benchmark datasets.
Model DistPred iTransformer PatchTST SCINet TimesNet DLinear FEDformer Autoformer Informer
Input Length 96 96 336 168 96 336 96 96 96
Output CRPS QICE PICP MSE MAE MSE MAE MSE MAE MSE MAE MSE MAE MSE MAE MSE MAE MSE MAE MSE MAE
ETT 96 0.248 9.53 48.50 0.288 0.334 0.297 0.349 0.302 0.348 0.707 0.621 0.340 0.374 0.333 0.387 0.358 0.397 0.346 0.388 3.755 1.525
192 0.277 9.22 51.06 0.348 0.371 0.380 0.400 0.388 0.400 0.860 0.689 0.402 0.414 0.477 0.476 0.429 0.439 0.456 0.452 5.602 1.931
336 0.298 8.91 53.28 0.392 0.402 0.428 0.432 0.426 0.433 1.000 0.744 0.452 0.452 0.594 0.541 0.496 0.487 0.482 0.486 4.721 1.835
720 0.322 8.23 57.11 0.437 0.436 0.427 0.445 0.431 0.446 1.249 0.838 0.462 0.468 0.831 0.657 0.463 0.474 0.515 0.511 3.647 1.625
Avg 0.286 8.97 52.45 0.366 0.386 0.383 0.407 0.387 0.407 0.954 0.723 0.414 0.427 0.559 0.515 0.437 0.449 0.450 0.459 4.431 1.729
Traffic 96 0.193 12.75 34.85 0.391 0.251 0.395 0.268 0.544 0.359 0.788 0.499 0.593 0.321 0.650 0.396 0.587 0.366 0.613 0.388 0.719 0.391
192 0.197 12.81 34.06 0.416 0.269 0.417 0.276 0.540 0.354 0.530 0.505 0.617 0.336 0.598 0.370 0.604 0.373 0.616 0.382 0.696 0.379
336 0.202 12.90 34.03 0.427 0.275 0.433 0.283 0.551 0.358 0.558 0.508 0.629 0.336 0.605 0.373 0.621 0.383 0.622 0.337 0.777 0.420
720 0.218 - - 0.467 0.297 0.467 0.302 0.586 0.375 0.841 0.523 0.640 0.350 0.645 0.394 0.626 0.382 0.660 0.408 0.864 0.472
Avg 0.203 - - 0.425 0.273 0.428 0.282 0.555 0.362 0.804 0.509 0.620 0.336 0.625 0.383 0.610 0.376 0.628 0.379 0.764 0.416
Electricity 96 0.167 12.16 34.11 0.138 0.231 0.148 0.240 0.195 0.285 0.247 0.345 0.168 0.272 0.197 0.282 0.193 0.308 0.201 0.317 0.274 0.368
192 0.178 11.62 38.85 0.155 0.246 0.162 0.253 0.199 0.289 0.257 0.355 0.184 0.289 0.196 0.285 0.201 0.315 0.222 0.334 0.296 0.386
336 0.188 11.14 41.27 0.169 0.264 0.178 0.269 0.215 0.305 0.269 0.369 0.198 0.300 0.209 0.301 0.214 0.329 0.231 0.338 0.300 0.394
720 0.209 10.79 42.83 0.207 0.298 0.225 0.317 0.256 0.337 0.299 0.390 0.220 0.320 0.245 0.333 0.246 0.355 0.254 0.361 0.373 0.439
Avg 0.186 11.45 39.27 0.167 0.260 0.178 0.270 0.216 0.304 0.268 0.365 0.192 0.295 0.212 0.300 0.214 0.327 0.227 0.338 0.311 0.397
Weather 96 0.145 12.05 29.96 0.152 0.192 0.174 0.214 0.177 0.218 0.221 0.306 0.172 0.220 0.196 0.255 0.217 0.296 0.266 0.336 0.300 0.384
192 0.185 10.607 39.85 0.210 0.247 0.221 0.254 0.225 0.259 0.261 0.340 0.219 0.261 0.237 0.296 0.276 0.336 0.307 0.367 0.598 0.544
336 0.216 9.857 44.84 0.263 0.286 0.278 0.296 0.278 0.297 0.309 0.378 0.280 0.306 0.283 0.335 0.339 0.380 0.359 0.395 0.578 0.523
720 0.263 9.31 47.83 0.362 0.349 0.358 0.349 0.354 0.348 0.377 0.427 0.365 0.359 0.345 0.381 0.403 0.428 0.419 0.428 1.059 0.741
Avg 0.202 10.46 40.62 0.247 0.269 0.258 0.279 0.259 0.281 0.292 0.363 0.259 0.287 0.265 0.317 0.309 0.360 0.338 0.382 0.634 0.548
Solar 96 0.152 6.36 65.75 0.205 0.227 0.203 0.237 0.234 0.286 0.237 0.344 0.250 0.292 0.290 0.378 0.242 0.342 0.884 0.711 0.236 0.259
192 0.165 8.39 53.34 0.236 0.251 0.233 0.261 0.267 0.310 0.280 0.380 0.296 0.318 0.320 0.398 0.285 0.380 0.834 0.692 0.217 0.269
336 0.171 7.82 55.37 0.256 0.264 0.248 0.273 0.290 0.315 0.304 0.389 0.319 0.330 0.353 0.415 0.282 0.376 0.941 0.723 0.249 0.283
720 0.187 8.79 55.15 0.273 0.278 0.249 0.275 0.289 0.317 0.308 0.388 0.338 0.337 0.356 0.413 0.357 0.427 0.882 0.717 0.241 0.317
Avg 0.169 7.84 57.40 0.243 0.255 0.233 0.262 0.270 0.307 0.282 0.375 0.301 0.319 0.330 0.401 0.291 0.381 0.885 0.711 0.235 0.280
PEMS 12 0.120 9.01 52.99 0.064 0.166 0.071 0.174 0.099 0.216 0.066 0.172 0.085 0.192 0.122 0.243 0.126 0.251 0.272 0.385 0.126 0.233
24 0.141 9.05 52.01 0.087 0.192 0.093 0.201 0.142 0.259 0.085 0.198 0.118 0.223 0.201 0.317 0.149 0.275 0.334 0.440 0.139 0.250
36 0.170 8.41 57.39 0.124 0.233 0.125 0.236 0.211 0.319 0.127 0.238 0.155 0.260 0.333 0.425 0.227 0.348 1.032 0.782 0.186 0.289
48 0.181 9.37 49.55 0.140 0.245 0.160 0.270 0.269 0.370 0.178 0.287 0.228 0.317 0.457 0.515 0.348 0.434 1.031 0.796 0.233 0.323
Avg 0.153 8.96 52.99 0.104 0.209 0.113 0.221 0.180 0.291 0.114 0.224 0.147 0.248 0.278 0.375 0.213 0.327 0.667 0.601 0.171 0.274
# Top 1 - - - 22 28 5 1 0 1 1 0 0 0 1 0 0 0 0 0 2 0
3.6 Visualization of the predictive distribution
The predictive distribution of DistPred is visualized in Figure 7. It can be observed that DistPred gives an ensemble of predictions (only top 10 are presented in the left subplot). Given all predictive ensemble values, the model can estimate the distribution of the response variable. Therefore, we can calculate confidence intervals at different levels, as shwon in the right subplot of Figure 7.
Figure 7: Visualization of the prediction results and the confidence intervals.
4 Conclusion
In this paper, we propose a novel method called DistPred, which is a distribution-free probabilistic inference method for regression and forecasting tasks. We transform proper scoring rules that measure the discrepancy between the predicted distribution and the target distribution into a differentiable discrete form and use it as a loss function to train the model end-to-end. This allows the model to sample numerous samples in a single forward pass to estimate the potential distribution of the response variable. We also propose a hybrid method called DistPred-MCD, which combines DistPred with MC Dropout to further improve the performance of uncertain quantization. Experimental results demonstrate that DistPred outperforms existing methods, often by a considerable margin. We also extend time series forecasting from point estimation to distribution prediction and achieve state-of-the-art performance on multivariate and univariate time series forecasting tasks. In the future, we plan to extend DistPred to other tasks, such as classification and reinforcement learning.
Appendix A Related work
In supervised learning contexts, the endeavor to characterize the conditional distribution $p(y|x)$ beyond merely the conditional mean $\mathop{\mathbb{E}}[y|x]$ via deep neural networks has been a focal point of existing research efforts. These endeavors primarily concentrate on quantifying predictive uncertainty, with several approaches having been proposed. Bayesian neural networks (BNNs) represent one such approach, aiming to capture such uncertainty by positing distributions over network parameters, thereby encapsulating the model’s plausibility given the available data (Blundell et al., 2015; Hernández-Lobato and Adams, 2015; Gal and Ghahramani, 2016; Kingma et al., 2015; Tomczak et al., 2021). Another avenue is represented by Kendall and Gal (2017), which not only addresses uncertainties in model parameters but also incorporates an additive noise term into the neural network output to encompass uncertainties in model outputs. In parallel, ensemble-based methodologies (Lakshminarayanan et al., 2017; Liu et al., 2022b) have emerged to address predictive uncertainty. These methods involve amalgamating multiple neural networks with stochastic outputs, thereby providing a comprehensive approach to uncertainty quantification. Furthermore, the neural processes’ family (Garnelo et al., 2018b, a; Kim et al., 2019; Gordon et al., 2020) has introduced a suite of models tailored to capturing predictive uncertainty in a manner that extends beyond the distribution of available data, particularly tailored for few-shot learning scenarios.
The aforementioned models have predominantly operated under the assumption of a parametric form in $p(y|x)$, typically adopting a Gaussian distribution or a mixture of Gaussians. They optimize network parameters by minimizing the negative log-likelihood of a Gaussian objective function. In contrast, deep generative models are renowned for their capacity to model implicit distributions without relying on parametric distributional assumptions. However, only a sparse number of works have ventured into leveraging this capability to address regression tasks. GAN-based models, as introduced by Zhou et al. (2021b) and Liu et al. (2021), have emerged as one such endeavor, focusing on conditional density estimation and predictive uncertainty quantification. Additionally, Han et al. (2022) have proposed a diffusion-based model tailored for conditional density estimation. Nevertheless, it is imperative to note that these models entail protracted training processes and computationally demanding inference procedures.
Appendix B Dataset
B.1 Commonly used TS datasets
Table 4: Details of the seven TS datasets.
Dataset length features frequency
ETTh1 17,420 7 1h
ETTh2 17,420 7 1h
ETTm1 69,680 7 15m
ETTm2 69,680 7 15m
Electricity 26,304 321 1h
Exchange 7,588 8 1d
Traffic 17,544 862 1h
Weather 52,696 21 10m
Solar 52,560 137 10m
PEMS 26,208 358 5m
Illness 966 7 7d
The information of the experiment datasets used in this paper are summarized as follows: (1) Electricity Transformer Temperature (ETT) dataset Zhou et al. (2021a), which contains the data collected from two electricity transformers in two separated counties in China, including the load and the oil temperature recorded every 15 minutes (ETTm) or 1 hour (ETTh) between July 2016 and July 2018. (2) Electricity (ECL) dataset  [^1] https://archive.ics.uci.edu/ml/datasets/ElectricityLoadDiagrams20112014 collects the hourly electricity consumption of 321 clients (each column) from 2012 to 2014. (3) Exchange Lai et al. (2018) records the current exchange of 8 different countries from 1990 to 2016. (4) Traffic dataset  [^2] http://pems.dot.ca.gov records the occupation rate of freeway system across State of California measured by 861 sensors. (5) Weather dataset  [^3] https://www.bgc-jena.mpg.de/wetter records every 10 minutes for 21 meteorological indicators in Germany throughout 2020. (6) Solar-Energy Lai et al. (2018) documents the solar power generation of 137 photovoltaic (PV) facilities in the year 2006, with data collected at 10-minute intervals. (7) The PEMS dataset Liu et al. (2022a) comprises publicly available traffic network data from California, collected within 5-minute intervals and encompassing 358 attributes. (8) Illness (ILI) dataset  [^4] https://gis.cdc.gov/grasp/fluview/fluportaldashboard.html describes the influenza-like illness patients in the United States between 2002 and 2021, which records the ratio of patients seen with illness and the total number of the patients. The detailed statistics information of the datasets is shown in Table 4.
The dataset information in terms of their size and number of features is summarized in Table 5.
Table 5: Dataset size ($N$ observations, $P$ features) of UCI regression tasks.
Dataset Boston Concrete Energy Kin8nm Naval Power Protein Wine Yacht Year
$(N,P)$ $(506,13)$ $(1030,8)$ $(768,8)$ $(8192,8)$ $(11,934,16)$ $(9568,4)$ $(45,730,9)$ $(1599,11)$ $(308,6)$ $(515,345,90)$
Table 6: Batch size settings of UCI regression tasks across different models.
PBP MC Dropout Deep Ensembles GCDS CARD (ours)
Boston $32$ $32$ $32$ $32$ $32$
Concrete $32$ $32$ $32$ $32$ $32$
Energy $32$ $32$ $32$ $32$ $32$
Kin8nm $64$ $32$ $64$ $64$ $64$
Naval $64$ $32$ $64$ $64$ $64$
Power $64$ $64$ $64$ $64$ $64$
Protein $100$ $256$ $100$ $256$ $256$
Wine $32$ $32$ $32$ $32$ $32$
Yacht $32$ $32$ $32$ $32$ $32$
Year $256$ $256$ $100$ $256$ $256$
B.2 Distribution free vs. distribution related
If we assume that the response variable follows a continuous distribution, as done in (Salinas et al., 2020) where $\hat{y}$ is assumed to be followed Gaussian distribution, we can provide an analytical formula for the Gaussian likelihood. Specifically, if $\hat{y}\sim\mathcal{N}(\mu,\sigma)$, then we can parametrize the Gaussian likelihood using its mean and standard deviation,
$\mathcal{L}_{G}(\hat{y}|\mu,\sigma)=(2\pi\sigma^{2})^{-\frac{1}{2}}exp(-(\hat{y}-\mu)^{2}/(2\sigma^{2})).$ (13)
Then, we can train DistPred utilizing Equation 13 as the loss function. Currently, DistPred exclusively provides the mean and variance of the response variable.
Table 7: Comparison of DistPred trained with Gaussia distribution.
Ettm2 Traffic Exchange Weather
CRPS MSE MAE CRPS MSE MAE CRPS MSE MAE CRPS MSE MAE
DistPred 0.189 0.12 0.256 0.161 0.137 0.216 0.343 0.426 0.456 0.023 0.002 0.028
DistPred-G 4391.80 7.61 1.73 2329.05 276.51 6.35 177.92 16.52 1.72 229.58 2.27 0.17
Appendix C Implementation details
The model undergoes training utilizing the ADAM optimizer Kingma and Ba (2015) and minimizing the Mean Squared Error (MSE) loss function. The training process is halted prematurely, typically within 10 epochs. The DistPred architecture solely comprises the embedding layer and backbone architecture, devoid of any additional introduced hyperparameters. During model validation, two evaluation metrics are employed: CRPS, QICE, PICP, MSE and MAE. Given the potential competitive relationship between the two indicators, MSE and MAE, we use the average of the two ($\frac{MSE+MAE}{2}$) to evaluate the overall performance of the model.
Appendix D Univariate time series forecasting
The full results for univariate TS forecasting are presented in Table 8. As other models, e.g., iTransformer Liu et al. (2023) and PatchTST Nie et al. (2022)do not offer performance information for all prediction lengths, we compare our method with those that provide comprehensive performance analysis, including FEDformer Zhou et al. (2022), Autoformer Wu et al. (2021), Informer Zhou et al. (2021a), LogTrans Li et al. (2019) and Reformer Kitaev et al. (2020). This reaffirms the effectiveness of DistPred.
Table 8: Univariate time series forecasting results on benchmark datasets.
Model DistPred-96 FEDformer-96 Autoformer-96 Informer-96 LogTrans-96 Reformer-96
Length CRPS QICE PICP MSE MAE MSE MAE MSE MAE MSE MAE MSE MAE MSE MAE
ETTh1 96 0.134 6.341 96.936 0.057 0.182 0.079 0.215 0.071 0.206 0.193 0.377 0.283 0.468 0.532 0.569
192 0.152 3.766 97.479 0.073 0.207 0.104 0.245 0.114 0.262 0.217 0.395 0.234 0.409 0.568 0.575
336 0.162 3.762 98.071 0.080 0.221 0.119 0.270 0.107 0.258 0.202 0.381 0.386 0.546 0.635 0.589
720 0.164 4.176 88.437 0.082 0.226 0.142 0.299 0.126 0.283 0.183 0.355 0.475 0.628 0.762 0.666
Avg 0.153 4.505 95.235 0.073 0.209 0.111 0.257 0.105 0.252 0.199 0.377 0.345 0.513 0.624 0.600
ETTh2 96 0.203 7.511 81.780 0.129 0.276 0.128 0.271 0.153 0.306 0.213 0.373 0.217 0.379 1.411 0.838
192 0.240 4.164 80.855 0.176 0.327 0.185 0.33 0.204 0.351 0.227 0.387 0.281 0.429 5.658 1.671
336 0.281 5.482 87.817 0.234 0.348 0.231 0.378 0.246 0.389 0.242 0.401 0.293 0.437 4.777 1.582
720 0.274 5.820 89.181 0.219 0.379 0.278 0.42 0.268 0.409 0.291 0.439 0.218 0.387 2.042 1.039
Avg 0.250 5.744 84.908 0.190 0.341 0.206 0.350 0.218 0.364 0.243 0.400 0.252 0.408 3.472 1.283
ETTm1 96 0.095 6.492 88.992 0.029 0.126 0.033 0.140 0.056 0.183 0.109 0.277 0.049 0.171 0.296 0.355
192 0.117 4.496 92.619 0.044 0.158 0.058 0.186 0.081 0.216 0.151 0.310 0.157 0.317 0.429 0.474
336 0.137 5.488 92.264 0.058 0.186 0.084 0.231 0.076 0.218 0.427 0.591 0.289 0.459 0.585 0.583
720 0.163 3.541 95.713 0.080 0.218 0.102 0.250 0.110 0.267 0.438 0.586 0.430 0.579 0.782 0.73
Avg 0.128 5.004 92.397 0.053 0.172 0.069 0.202 0.081 0.221 0.281 0.441 0.231 0.382 0.523 0.536
ETTm2 96 0.133 9.595 69.279 0.064 0.180 0.063 0.189 0.065 0.189 0.08 0.217 0.075 0.208 0.077 0.214
192 0.173 7.392 75.604 0.099 0.233 0.110 0.252 0.118 0.256 0.112 0.259 0.129 0.275 0.138 0.290
336 0.203 8.833 82.711 0.133 0.277 0.147 0.301 0.154 0.305 0.166 0.314 0.154 0.302 0.160 0.313
720 0.248 5.720 90.470 0.185 0.333 0.219 0.368 0.182 0.335 0.228 0.380 0.160 0.322 0.168 0.334
Avg 0.189 7.885 79.516 0.120 0.256 0.135 0.278 0.130 0.271 0.147 0.293 0.130 0.277 0.136 0.288
Traffic 96 0.155 4.245 25.012 0.132 0.209 0.170 0.263 0.246 0.346 0.257 0.353 0.226 0.317 0.313 0.383
192 0.158 3.596 25.458 0.136 0.213 0.173 0.265 0.266 0.37 0.299 0.376 0.314 0.408 0.386 0.453
336 0.171 3.458 25.146 0.134 0.213 0.178 0.266 0.263 0.371 0.312 0.387 0.387 0.453 0.423 0.468
720 0.135 3.434 24.926 0.146 0.228 0.187 0.286 0.269 0.372 0.366 0.436 0.437 0.491 0.378 0.433
Avg 0.161 3.683 25.136 0.137 0.216 0.177 0.220 0.261 0.365 0.309 0.388 0.341 0.417 0.375 0.434
Electricity 96 0.266 5.619 51.949 0.257 0.366 0.262 0.378 0.341 0.438 0.258 0.367 0.288 0.393 0.275 0.379
192 0.276 5.018 65.785 0.284 0.376 0.316 0.410 0.345 0.428 0.285 0.388 0.432 0.483 0.304 0.402
336 0.317 4.947 62.019 0.400 0.439 0.361 0.445 0.406 0.470 0.336 0.423 0.430 0.483 0.37 0.448
720 0.335 5.191 67.705 0.409 0.460 0.448 0.501 0.565 0.581 0.607 0.599 0.491 0.531 0.46 0.511
Avg 0.299 5.191 61.865 0.337 0.410 0.347 0.434 0.414 0.479 0.372 0.444 0.410 0.473 0.352 0.435
Weather 96 0.019 10.732 61.824 0.0012 0.024 0.0035 0.046 0.0110 0.081 0.004 0.044 0.0046 0.052 0.012 0.087
192 0.022 10.463 55.581 0.0013 0.027 0.0054 0.059 0.0075 0.067 0.002 0.040 0.006 0.060 0.0098 0.044
336 0.023 10.271 59.075 0.0021 0.024 0.008 0.072 0.0063 0.062 0.004 0.049 0.006 0.054 0.013 0.100
720 0.027 10.515 47.492 0.0023 0.033 0.015 0.091 0.0085 0.070 0.003 0.042 0.007 0.059 0.011 0.083
Avg 0.023 10.495 55.993 0.0023 0.028 0.008 0.067 0.0083 0.0700 0.0033 0.0438 0.0059 0.0563 0.0115 0.0785
Exchange 96 0.182 6.99 94.678 0.110 0.241 0.131 0.284 0.241 0.387 1.327 0.944 0.237 0.377 0.298 0.444
192 0.263 10.391 82.484 0.204 0.338 0.277 0.420 0.300 0.369 1.258 0.924 0.738 0.619 0.777 0.719
336 0.376 10.778 87.861 0.401 0.482 0.426 0.511 0.509 0.524 2.179 1.296 2.018 1.0700 1.833 1.128
720 0.552 9.526 81.032 0.991 0.763 1.162 0.832 1.260 0.867 1.28 0.953 2.405 1.175 1.203 0.956
Avg 0.343 9.421 86.514 0.426 0.456 0.499 0.512 0.578 0.537 1.511 1.029 1.350 0.810 1.028 0.812
$1^{\text{st}}$ Count - - - 34 37 3 1 0 0 1 1 2 1 0 0
Appendix E Limitations of DistPred
DistPred is not without its limitations. Firstly, the model is not designed to handle categorical data, and as such, it is not suitable for classification tasks. Secondly, the model is not designed to handle time series data with missing values. Thirdly, the model is not designed to handle multivariate time series data. Finally, the model is not designed to handle time series data with irregular time intervals.

### v2

@@
 Traditional regression and prediction tasks often only provide deterministic point estimates.
-To estimate the uncertainty or distribution information of the response variable, methods such as Bayesian inference, model ensembling, or MC Dropout are typically used.
-These methods either assume that the posterior distribution of samples follows a Gaussian process or require thousands of forward passes for sample generation.
+To estimate the distribution or uncertainty of the response variable, traditional methods either assume that the posterior distribution of samples follows a Gaussian process or require thousands of forward passes for sample generation.
 We propose a novel approach called DistPred for regression and forecasting tasks, which overcomes the limitations of existing methods while remaining simple and powerful.
@@
 Additionally, our method significantly improves computational efficiency.
-For example, compared to state-of-the-art models, DistPred has a 90x faster inference speed.
+For example, compared to state-of-the-art models, DistPred has a 180x faster inference speed.
 Experimental results can be reproduced through this Repository.
@@
 Moreover, this is particularly important in high-security AI application areas such as autonomous driving, risk estimation, and decision-making.
-In this paper, we consider the underlying distribution behind predicting the response variable because it reflects the confidence intervals of all levels.
-For example, based on this distribution, we can calculate confidence intervals at any level, coverage rate, and uncertainty quantification.
+In this paper, we consider the underlying distribution behind predicting the response variable because it reflects the confidence intervals at all levels.
+For example, based on this distribution, we can calculate confidence intervals, coverage rate, and uncertainty quantification at any level, as shown in Fig.
+1.
 Currently, predicting the distribution of the response variable poses a challenge because at a specific moment, the response variable can only take on a single deterministic value.
-This point can be viewed as a maximum likelihood sample from its underlying distribution, but it fails to reflect the overall state of the underlying distribution.
-Currently, the primary method used to address distribution prediction and uncertainty quantification in regression and prediction tasks is frequentist sampling.
-These methods involve sampling numerous samples by perturbing the explanatory variable or model to approximate the underlying distribution of the response variable.
-For instance, Bayesian neural networks (BNNs) simulate this uncertainty by assuming that their parameters follow a Gaussian distribution, thereby capturing the model’s uncertainty given the data (Blundell et al., 2015).
+This point can be viewed as a ‘representative’ sample from its underlying distribution, but it fails to represent the overall state of the underlying distribution.
+Figure 1: DistPred can provide $K$ predicted values $\hat{y}$ of the response variable $y$ given the predictor variable $x$ in a single forward process, denoted as $\mathbb{E}(\hat{Y}|x)$, where $\hat{Y}$ represents a maximum likelihood sample of $y$.
+Based on this sampling, the probability mass function (PMF) $P_{\theta}(\hat{y}|x)$, cumulative distribution function (CDF) $F_{\theta}(\hat{y}|x)$, and confidence curve $CC_{\theta}(\hat{y}|x)$ for the response variable $y$ can be computed, thereby yielding comprehensive statistical insights into $y$.
+For instance, this includes confidence intervals (CI) at any desired level, as well as p-values.
+Currently, several methods are employed to predict the underlying distribution of the response variable in regression and forecasting tasks.
+The most straightforward approach is to assume that the response variable follows a prior distribution and represents a specific statistic from that distribution.
+Specifically, several methods (Bishop, 1994; Greene, 2003; Salinas et al., 2020; He et al., 2020) transform distribution prediction and uncertainty quantification into predicting statistical variables such as mean and variance by assuming that the response variable follows a known continuous distribution.
+For instance, mixture density networks (Bishop, 1994) superimpose a specific distribution, typically Gaussian, weighted by designated parameters to fit the prior distribution.
+Heteroscedasticity regression (Greene, 2003) quantifies uncertainty by modeling the variability of residuals as a function of independent variables.
+These methods only predict statistical variables, which reduces the inference cost, but strong assumptions often fail to capture the true data distribution, resulting in inferior performance.
+Conformal prediction offers an alternative approach for distribution prediction (Vovk et al., 2017, 2018; Romano et al., 2019; Xu and Xie, 2021, 2023).
+The authors defined the random prediction system in (Vovk et al., 2017, 2018) and proposed a nonparametric prediction distribution method based on conformal assumptions.
+The authors integrated conformal prediction with quantile regression in (Romano et al., 2019; Xu and Xie, 2021, 2023) to construct prediction intervals for the response variable by training multiple bootstrap estimators.
+However, the application of conformal prediction to distribution prediction can be constrained by its reliance on the exchangeability assumption of residuals and the challenges in managing temporal dependencies (autocorrelation), potentially resulting in less reliable prediction intervals in non-i.i.d.
+data settings.
+Uncertainty quantification is a method that indirectly reflects the potential distribution of the response variable, which can be classified into two primary categories: epistemic uncertainty and aleatoric uncertainty Der Kiureghian and Ditlevsen (2009).
+Epistemic uncertainty refers to the uncertainty within the model itself.
+In contrast, aleatoric uncertainty pertains to the inherent randomness in the observations.
+Quantitative analysis involves generating numerous samples (e.g., using MCMC) by perturbing the explanatory variables or models, thereby approximating the underlying distribution.
+For instance, bayesian neural networks (BNNs) (Blundell et al., 2015; Immer et al., 2021; Daxberger et al., 2021) simulate this uncertainty by assuming that their parameters follow a predefined distribution, thereby capturing the model’s uncertainty given the data.
 Similarly, ensemble-based methods have been proposed to combine multiple deep models with random outputs to capture prediction uncertainty.
-MC Dropout (Gal and Ghahramani, 2016) demonstrates that enabling dropout during each testing process yields results akin to model ensembling.
-Additionally, models based on GANs and diffusion are introduced for conditional density estimation and prediction uncertainty quantification.
+MC Dropout (Gal and Ghahramani, 2016) shows that enabling dropout during each testing process yields results akin to model ensembling.
+Additionally, models based on GANs and diffusion have been introduced for conditional density estimation and prediction uncertainty quantification (Zhou et al., 2021b; Liu et al., 2021; Han et al., 2022).
 These models utilize noise during the generation or diffusion process to obtain different predicted values for estimating the uncertainty of the response variable.
 The common characteristic of these methods mentioned above is the requirement of $K$ forward passes to sample $K$ representative samples.
-For example, Bayesian framework-based methods require $K$ learnable parameter samples to be inferred to obtain $K$ representative samples; ensemble methods require $K$ models to jointly infer; MC Dropout requires $K$ forward passes with random dropout activation; generative models require $K$ forward or diffusion processes.
+For example, Bayesian framework-based methods require $K$ learnable parameter samples to be inferred in order to obtain $K$ representative samples; ensemble methods require $K$ models to jointly infer; MC Dropout requires $K$ forward passes with random dropout activations; generative models require $K$ forward or diffusion processes.
 However, the excessive forward passes result in significant computational overhead and slow speed, a drawback that becomes increasingly apparent for AI applications with high real-time requirements.
-Furthermore, several methods (Nix and Weigend, 1994; Salinas et al., 2020; He et al., 2020) transform distribution prediction and uncertainty quantification into predicting statistical variables such as mean and variance by assuming that the response variable follows a known continuous distribution.
-For instance, DeepAR (Salinas et al., 2020) assumes the response variable follows a Gaussian distribution, thereby directly utilizing GaussianNLLLoss (Nix and Weigend, 1994) to optimize its mean and variance.
-These methods only predict statistical variables, reducing the inference cost, but strong distribution assumptions often fail to capture the true data distribution, resulting in inferior performance.
-Additionally, there are methods that train models using quantiles as loss functions to obtain specific quantile intervals (QIs) of the response variable.
-For example, to predict the 5th and 95th percentile confidence intervals, one can directly predict the corresponding quantiles.
-MRQNN predicts specified quantiles using quantile loss.
-Similarly, TFT enables the model to directly output three quantiles: 0.1, 0.5, and 0.9.
-These methods cannot provide distribution information about the response variable but only offer fixed QIs, thus lacking flexibility in their predictions.
-Figure 1: DistPred can provide $K$ predicted values $\hat{y}$ of the response variable $y$ given the predictor variable $x$ in a single forward process, denoted as $\mathbb{E}(\hat{Y}|x)$, where $\hat{Y}$ represents a maximum likelihood sample of $y$.
-Based on this sampling, the probability mass distribution (PMD) $P_{\theta}(\hat{y}|x)$, cumulative distribution function (PDF) $F_{\theta}(\hat{y}|x)$, and confidence curve $CC_{\theta}(\hat{y}|x)$ for the response variable $y$ can be computed, thereby yielding comprehensive statistical insights into $y$.
-For instance, this includes confidence intervals (CI) at any desired level, as well as p-values.
 To address this issue, we propose a novel method called DistPred, which is a distribution-free probabilistic inference method for regression and forecasting tasks.
@@
 DistPred is orthogonal to other methods, enabling its combination with alternative approaches to enhance estimation performance.
-Further, we show that DistPred can provide comprehensive statistical insights into the response variable, including confidence intervals at any desired level, p-values, and other statistical information, as shown in Figure 1.
+Further, we show that DistPred can provide comprehensive statistical insights into the response variable, including confidence intervals at any desired level, p-values, and other statistical information, as shown in Fig.
+1.
 Experimental results show that DistPred outperforms existing methods in terms of both accuracy and computational efficiency.
-Specifically, DistPred has a 90x faster inference speed than state-of-the-art models.
+Specifically, DistPred has a 180x faster inference speed than state-of-the-art models.
+Figure 2: The workflow of DistPred. An ensemble of predictive variables $\hat{Y}$ is inferred in a forward pass and $S(\mathbb{E}(\hat{Y}|x),y)$ is utilized to train the learner end-to-end.
 2 Method
@@
 The subscript $i$ will be omitted if it does not cause ambiguity in the context.
-Our objective is to utilize a machine learning model $M$ with parameters $\theta$ to predict the underlying distribution $P(y)$ of the response variable $y$ from $D$, aiming to acquire comprehensive statistical insights such as obtaining confidence intervals (CI) and quantifying uncertainty at any desired level.
-Direct prediction of distribution $P$ is not feasible, because:
+Our objective is to utilize a machine learning model $M$ with parameters $\theta$ to predict the underlying distribution $P(y|x)$ from $D$, aiming to acquire comprehensive statistical insights such as obtaining confidence intervals (CI) and quantifying uncertainty at any desired level.
+Direct prediction of distribution $P_{\theta}(\hat{y}|x)$ is not feasible because:
   1. 1.
-    Without distributional assumptions, we cannot give a valid representation of the PDF or CDF of the predicted distribution $P_{\theta}(\hat{y})$.
+    Without distributional assumptions, we cannot give a valid representation of the PDF or CDF for the response variable.
   2. 2.
-    We can only obtain a singular deterministic value for the response variable $y$, lacking access to its distributional information to guide model learning.
-Figure 2: Relationship between CDF and all predictive quantiles.
-To address the aforementioned issues, we contemplate employing all predictive quantiles $\hat{q}_{1},\hat{q}_{2},\cdots$, at levels $\alpha_{1},\alpha_{2},\cdots$ to specify the potential CDF $F_{\theta}(\hat{y})$ of the predictor variable $\hat{y}$.
-This is because if we know the cumulative distribution function of a random variable, we can find any quantile by setting $F(y)=q$.
-Conversely, if we have a complete set of quantiles, we can approximate or reconstruct the cumulative distribution function of the random variable.
-As shwon in Figure 2, quantiles provide discrete ‘snapshots’ of the distribution, while the CDF is a continuous, smooth version of these snapshots, offering a complete description of the cumulative probability from the minimum to the maximum value.
-Next, we will present that the full predictive quantiles outlined earlier serves as a proper approximation of the response variable’s CDF.
-And the full quantiles’ prediction can be translated into calculating the minimum expected score of the response variable and the predictive ensemble variables $\hat{Y}$, where $\hat{Y}=\{\hat{y}_{1},\cdots,\hat{y}_{K}\}$.
-Prior to delving into this analysis, we shall initially introduce the scoring rules utilized for assessing the propriety of predicted distributions.
-2.1 Utilize proper scoring rules as the loss function
-Scoring rules offer a concise measure for assessing probabilistic forecasts by assigning numerical scores according to the forecast distribution and predicted outcomes (Gneiting and Raftery, 2007; Jordan et al., 2017).
-Specifically, Let $\Omega$ denote the set of possible values of the quantity of interest, and let $\mathcal{P}$ denote a convex class of probability distributions on $\Omega$.
-A scoring rule is a function
+    We have only a deterministic target point, without access to its distribution information, which limits our ability to guide the model’s learning process.
+To address the aforementioned issues, we contemplate employing full predictive quantiles $\hat{q}_{1},\hat{q}_{2},\cdots,\hat{q}_{K}$ at levels $\alpha_{1},\alpha_{2},\cdots,\alpha_{K}$, $(K\rightarrow\infty)$, to specify the potential CDF $F_{\theta}(\hat{y})$ of the response variable $\hat{y}$.
+This is because if we know the cumulative distribution function of a random variable, we can find any quantile by
+$q_{\alpha}=\text{inf}\{y\in\mathbb{R}:F(y)\geq\alpha\}.$ (1)
+Conversely, if we have a complete set of quantiles, we can approximate or reconstruct the cumulative distribution function of the random variable
+$F(y)=\text{sup}\{\alpha\in[0,1]:q_{\alpha}\leq y\}.$ (2)
+Full quantiles provide discrete ‘snapshots’ of the distribution, while the CDF is a continuous, smooth version of these snapshots, offering a complete description of the cumulative probability from the minimum to the maximum value.
+2.1 Property of Full Quantiles
+We contemplate probabilistic predictions pertaining to a continuous quantity, manifested as full predictive quantiles $\hat{q}_{1},\cdots,\hat{q}_{K},(K\rightarrow\infty)$.
+For $P\in\mathcal{P}$, let $\hat{q}_{1},\cdots,\hat{q}_{K}$ denote the true $P$-quantiles at levels $\alpha_{1},\cdots,\alpha_{K}\in(0,1)$.
+Then, the expected score $S(q_{1},\cdots,q_{K};P)$ can be defined as
+$S(\hat{q}_{1},\cdots,\hat{q}_{K};P)=\int S(\hat{q}_{1},\cdots,\hat{q}_{K};y)\,\text{d}P(y).$ (3)
+The function $S$ in Eq.
+3 satisfies the scoring rule, which offers a concise measure for assessing probabilistic forecasts by assigning numerical scores according to the forecast distribution and predicted outcomes (Gneiting and Raftery, 2007; Jordan et al., 2017).
+Specifically, let $\Omega$ denote the set of possible values of the quantity of interest, and let $\mathcal{P}$ denote a convex class of probability distributions on $\Omega$, the scoring rule is a function
 $S:\Omega\times\mathcal{P}\rightarrow\mathbb{R}\cup\{\infty\}$ (4)
@@
 A scoring rule is termed strictly proper when equality is achieved only when $P=Q$.
-Proper scoring rules (PSR) are essential for comparative evaluation, particularly in ranking forecasts.
-In practice, the forecaster with the lowest average score across multiple forecast cases typically demonstrates the best predictive performance.
-Proper scoring rules incentivize forecasters to accurately report their perception of the true distribution in this scenario.
-Therefore, PSR provide attractive loss and utility functions that can be tailored to a regression or forecast problem.
-To estimate $\theta$, we might measure the goodness-of-fit by the mean score
-$S_{n}(\theta)=\frac{1}{N}\sum_{i=1}^{N}S(P_{\theta}(\hat{y}_{i}),y_{i}).$ (3)
-Let $\theta^{*}$ denotes the true parameter value, then asymptotic arguments indicate that $\text{argmin}_{\theta}S_{n}(\theta)\rightarrow\theta^{*}$ as $n\rightarrow\infty$.
-This suggests a general approach to transform PSR into loss functions for training models, which implicitly minimizes the divergence between predictive and true distributions.
-2.2 Property of full predictive quantiles
-We contemplate probabilistic predictions pertaining to a continuous quantity, manifested as full predictive quantiles $\hat{q}_{1},\cdots,\hat{q}_{K}$.
-For $P\in\mathcal{P}$, let $q_{1},\cdots,q_{K}$ denote the true $P$-quantiles at levels $\alpha_{1},\cdots,\alpha_{K}\in(0,1)$.
-Then, the expected score $S(q_{1},\cdots,q_{K};P)$ can be defined as
-$S(\hat{q}_{1},\cdots,\hat{q}_{K};P)=\int S(\hat{q}_{1},\cdots,\hat{q}_{K};y)\,\text{d}P(y).$ (4)
-Further, a scoring rule S is proper if $S(q_{1},\cdots,q_{K};P)\geq S(\hat{q}_{1},\cdots,\hat{q}_{K};P)$.
-Based on this definition, we assuming that $s_{k},i\in[1,\cdots K]$ is nondecreasing and $h$ is arbitrary, then the scoring rule
-$S(\hat{q}_{1},\cdots,\hat{q}_{K};P)=\sum_{k=1}^{K}\left(\alpha_{i}s_{k}(\hat{q}_{k})+(s_{k}(y)-s_{k}(\hat{q_{k}})\mathbbm{1}\{y\leq\hat{q}_{k}\})\right)$ (5)
+Proper scoring rules are essential for comparative evaluation, especially in ranking forecasts.
+Based on this definition, we can get the scoring rule $S$ is proper when $S(q_{1},\cdots,q_{K};P)\geq S(\hat{q}_{1},\cdots,\hat{q}_{K};P)$.
+Further, assuming that $s_{k}$ is a non-decreasing probability measure of $\mathcal{P}$, then the scoring rule
+$S(\hat{q}_{1},\cdots,\hat{q}_{K};P)=\sum_{k=1}^{K}\left(\alpha_{k}s_{k}(\hat{q}_{k})+(s_{k}(y)-s_{k}(\hat{q_{k}})\mathbb{I}\{y\leq\hat{q}_{k}\})\right)$ (6)
 is proper for predicting the quantiles at levels $\alpha_{1},\cdots,\alpha_{K}$ when $K\rightarrow\infty$ (Gneiting and Raftery, 2007; Schervish et al., 2012).
-$\mathbbm{1}\{y\leq\hat{q}_{k}\}$ denotes the indicator function which is one if $y\leq\hat{q}_{k}$ and zero otherwise.
-Equation 5 shows that full predictive quantiles is proper.
-In essence, delineating a predictive CDF is equivalent to specifying all predictive quantiles.
+$\mathbb{I}\{y\leq\hat{q}_{k}\}$ denotes the indicator function which is $1$ if $y\leq\hat{q}_{k}$ and 00 otherwise.
+2.2 Predicting CDF by Full Quantiles
+Eq.
+6 shows that full predictive quantiles are proper.
 Consequently, we can formulate scoring rules for the predictive distribution based on the scoring rules for the quantiles.
-Specifically, let $S_{\alpha}$ denotes a proper scoring rule for the quantile at level $\alpha$, then the scoring rule
-$S(F,y)=\int_{0}^{1}S_{\alpha}(F^{-1}(\alpha);y)\,\text{d}\alpha=\int_{-\infty}^{\infty}S(F(\hat{y}),\mathbbm{1}\{y\leq\hat{y}\})\,\text{d}\hat{y}$ (6)
-is proper. Here, we can find that the right of Equation 6 corresponds to the CRPS in which $S$ is the quadratic or Brier score, which is defined as
-$C(F,y)=\int_{-\infty}^{\infty}(F(\hat{y})-\mathbbm{1}\{y\leq\hat{y}\})^{2}\,\text{d}\hat{y}.$ (7)
-If the first moment of $F$ is finite, the CRPS can be written as
+Specifically, let $S_{\alpha}$ denote a proper scoring rule for the quantile at level $\alpha$, then the scoring rule
+$S(F,y)=\int_{0}^{1}S_{\alpha}(F^{-1}(\alpha);y)\,\text{d}\alpha=\int_{-\infty}^{\infty}S(F(\hat{y}),\mathbb{I}\{y\leq\hat{y}\})\,\text{d}\hat{y}$ (7)
+is proper.
+Here, we establish a relationship between the full quantiles of the response variable and its CDF.
+However, the CDF here is in continuous form and cannot be directly observed.
+Therefore, we need to convert it into a discrete form, as the number of quantiles $K$ is always finite in practice.
+It is worth noting that Eq.
+7 corresponds to the continuous ranked probability score (CRPS) in which $S$ is the quadratic or Brier score, defined as
+$C(F,y)=\int_{-\infty}^{\infty}(F(\hat{y})-\mathbb{I}\{y\leq\hat{y}\})^{2}\,\text{d}\hat{y}.$ (8)
+When the first moment of $F$ is finite, the CRPS can be written as
 $C(F,y)=\mathbb{E}_{F}[|\hat{Y}-y|]-\frac{1}{2}\mathbb{E}_{FF}[|\hat{Y}-\hat{Y}^{\prime}|],$ (9)
-where $\hat{Y}$ and $\hat{Y}^{\prime}$ denote independent predictive variables with distribution $F$.
-Figure 3: The workflow of DistPred. An ensemble of predictive variables $\hat{Y}$ is inferred in a forward pass and PSR $S(\mathbb{E}(\hat{Y}|x),y)$ is utilized to train the learner end-to-end.
+where $\hat{Y}$ and $\hat{Y}^{\prime}$ denote independent predictive ensemble variables with distribution $F$.
+Eq.
+9 incentivizes forecasters to accurately report their perception of the true distribution in this scenario.
+Therefore, it provides attractive measures and utility functions that can be tailored to a regression or forecast problem.
+To estimate $\theta$, we need to measure the goodness-of-fit by the mean score
+$C_{N}(\theta)=\frac{1}{N}\sum_{i=1}^{N}C(F_{\theta}(\hat{y}_{i}),y_{i}).$ (10)
+Let $\theta^{*}$ denote the true parameter value, then asymptotic arguments indicate that $\text{argmin}_{\theta}C_{N}(\theta)\rightarrow\theta^{*}$ as $N\rightarrow\infty$.
+As shown in Fig.
+2, the workflow suggests a general approach to transforming proper scoring rules into loss functions for training models, which implicitly minimizes the discrepancy between predictive and true distributions.
 2.3 End-to-End Ensemble Inference
-Figure 4: DistPred’s architecture.
-Based on the analysis provided above, it is evident that predicting the full quantiles is equivalent to minimizing Equation 8 w.r.t $\mathbb{E}(\hat{Y}|y)$.
-Hence, as the workflow shown in Figure 3, we can develop a model $M$ with parameters $\theta$ that infers an ensemble of predictive variables $\hat{Y}$ in a forward pass and utilize Equation 8 to train it end-to-end.
+Figure 3: DistPred’s architecture. A model $M$ with parameters $\theta$ takes one input variable $x$ and outputs an ensemble of $K$ response variable $\hat{y}_{1},\cdots,\hat{y}_{K}$.
+Based on the analysis provided above, it is evident that predicting the full quantiles is equivalent to minimizing Eq.
+9 w.r.t $\mathbb{E}(\hat{Y}|y)$.
+Hence, as illustrated in the architecture depicted in Fig.
+3, we can develop a model $M$ with parameters $\theta$ that infers an ensemble of predictive variables $\hat{Y}=\{\hat{y}_{1},\cdots,\hat{y}_{K}\}$ in a forward pass and utilize Eq.
+10 to train it end-to-end.
 This allows the model to sample numerous samples in a single forward pass to estimate the empirical CDF $\hat{F}$ by the predictive ensemble variables
-$C(\hat{F},y)=\frac{1}{K}\sum_{k=1}^{K}|\hat{y}_{k}-y|-\frac{1}{2K^{2}}\sum_{k=1}^{K}\sum_{j=1}^{K}|\hat{y}_{k}-\hat{y}^{\prime}_{j}|.$ (9)
-It’s worth noting that Equation 9 is a differentiable discrete form w.r.t $\hat{Y}$ and $\hat{Y^{\prime}}$ that strictly satisfies PSR.
-However, implementations of Equation 9 exhibit inefficiency due to their computational complexity of $O(K^{2})$.
-This can be enhanced by employing representations based on the generalized quantile function (Laio and Tamea, 2007) and the sorted predictive ensemble variables
-$C(\hat{F},y)=\frac{1}{2K^{2}}\sum_{k=1}^{K}(\overrightarrow{y_{k}}-y)(k\mathbbm{1}\{y\leq\overrightarrow{y_{k}}\}-i+\frac{1}{2}).$ (10)
-Since a sorting operation is involved, the computational complexity of Equation 10 is $O(KlogK)$.
-To conserve memory, we also suggest utilizing Equation 10 as the loss function, as predicting ensemble variables in long-term forecasting tasks may lead to out-of-memory issues on GPUs.
+$C(\hat{F},y)=\frac{1}{K}\sum_{k=1}^{K}|\hat{y}_{k}-y|-\frac{1}{2K^{2}}\sum_{k=1}^{K}\sum_{j=1}^{K}|\hat{y}_{k}-\hat{y}_{j}|.$ (11)
+It’s worth noting that Eq.
+11 is a differentiable discrete form w.r.t $\hat{Y}$ and $\hat{Y^{\prime}}$ that strictly satisfies proper scoring rules.
+However, implementations of Eq.
+11 exhibit inefficiency due to their storage complexity of $\mathcal{O}(K^{2})$.
+This can be enhanced by using algebraically equivalent representations based on the generalized quantile function (Laio and Tamea, 2007) and the sorted predictive ensemble variables $\overrightarrow{y_{k}}$,
+$C(\hat{F},y)=\frac{1}{2K^{2}}\sum_{k=1}^{K}(\overrightarrow{y_{k}}-y)(k\mathbb{I}\{y\leq\overrightarrow{y_{k}}\}-k+\frac{1}{2}).$ (12)
+Since a sorting operation is involved, the storage complexity of Eq.
+12 is $\mathcal{O}(KlogK)$.
+However, since Eq.
+12 renders the objective $C(\hat{F},y)$ non-differentiable, we employ probability-weighted moment estimation form to approximate it, as
+$C(\hat{F},y)=\frac{1}{K}\sum_{k=1}^{K}|\hat{y}_{k}-y|+\frac{1}{K}\sum_{k=1}^{K}\hat{y}_{k}-\frac{2}{K(K-1)}\sum_{k=1}^{K}\hat{y}_{k}(k-1).$ (13)
+Actually, according to the definition of L-Moments, we can prove that the last term of Eq.
+11 can be decomposed into the 1-th and 0-st order L-moments of Eq.
+13, i.e., Eq.
+13 is an unbiased estimate of Eq.
+11 when it has a finite first L-moment.
+To conserve memory, we suggest utilizing Eq.
+13, which has a storage complexity of $\mathcal{O}(K)$, as the loss function, since predicting ensemble variables in long-term forecasting tasks may lead to out-of-memory issues on GPUs.
 2.4 Incorporate Alternative Methodologies
 DistPred is orthogonal to other methods, enabling its combination with alternative approaches to enhance estimation performance.
-Here, with a focus on computational efficiency and memory conservation, we opt to integrate MC Dropout with DistPred, thereby denoting the amalgamation as DistPred-MCD.
-In our experiments, we observed that DistPred-MCD can further enhance uncertainty quantification performance, albeit with a marginal increase in computational effort.
+Here, with a focus on computational efficiency and memory conservation, we opt to integrate MC Dropout with DistPred, thereby denoting the amalgamation as DistPred+MCD.
+In our experiments, we observed that using DistPred+MCD can further enhance uncertainty quantification performance, albeit with a marginal increase in computational effort.
+Figure 4: Scatter plot of DistPred’s regression results on $8$ toy examples.
 3 Experiments
@@
 3.1 PICP and QICE Metrics
-In recent years, BNNs (Hernández-Lobato and Adams, 2015; Gal and Ghahramani, 2016) have emerged as a class of models that aims at estimating the uncertainty providing a more complete picture of $p(y|x)$.
-In (Han et al., 2022), it is contended that the metric utilized in BNNs to assess uncertainty estimates, namely the negative log-likelihood (NLL), is computed based on Gaussian densities.
+The metric utilized in BNNs to assess uncertainty estimates, namely the negative log-likelihood (NLL), is computed based on Gaussian prior.
 This assumption implies that they consider both the conditional distribution $p(y|x=x^{\prime})$ for all $x^{\prime}$ are Gaussian.
 However, this assumption is very difficult to verify for real-world datasets.
-We follow CARD (Han et al., 2022) and use the following two metrics, both of which are designed to empirically evaluate the degree of similarity between learned and true conditional distributions:
+We follow (Han et al., 2022) and use the following two metrics, both of which are designed to empirically evaluate the degree of similarity between learned and true conditional distributions:
   • PICP (Prediction Interval Coverage Probability) (Yao et al., 2019) is a metric that measures the proportion of true labels that fall within the prediction interval.
@@
 The PICP is calculated as
-$PICP:=\frac{1}{N}\sum_{n=1}^{N}\mathbbm{1}\{\hat{y}_{n}\geq q_{\alpha/2}\}\cdot\mathbbm{1}\{\hat{y}_{n}\leq q_{1-\alpha/2}\},$ (11)
+$PICP:=\frac{1}{N}\sum_{n=1}^{N}\mathbb{I}\{\hat{y}_{n}\geq q_{\alpha/2}\}\cdot\mathbb{I}\{\hat{y}_{n}\leq q_{1-\alpha/2}\},$ (14)
 where $q_{\alpha/2}$ and $q_{1-\alpha/2}$ represent the low and high percentiles, respectively, that we have selected for the predicted $\hat{y}$ outputs given the same $x$ input.
 This metric evaluates the proportion of accurate observations that lie within the percentile range of the generated $\hat{y}$ samples corresponding to each $x$ input.
-Within this study, we opt for the $2.5$th and $97.5$th percentiles, signifying that an optimal PICP value for the acquired model would ideally reach 95%.
+Within this study, we opt for the $2.5$th and $97.5$th percentiles, signifying that an optimal PICP value for the model should ideally reach 95%.
 However, a caveat of the PICP metric becomes apparent in the measurement of distribution differences.
-For instance, consider a situation where the 2.5th to 97.5th percentile range of the learned distribution encompasses the data falling between the 1st and 96th percentiles of the true distribution.
-Even with a sufficient number of samples, the resulting PICP value would still approach 95%.
-However, it becomes evident that a discrepancy exists between the learned distribution and the true distribution.
 Drawing from this reasoning, Han et al.
@@
 The definition of QICE entails computing the mean absolute error (MAE) between the proportion of true data encompassed by each quantile interval and the optimal proportion, which is 1/M for all intervals:
-$QICE:=\frac{1}{M}\sum_{m=1}^{M}|r_{m}-\frac{1}{M}|,\quad\text{where}\ \ r_{m}=\frac{1}{N}\sum_{n=1}^{N}\mathbbm{1}\{\hat{y}_{n}\geq q_{\alpha/2}\}\cdot\mathbbm{1}\{\hat{y}_{n}\leq q_{1-\alpha/2}\}.$ (12)
-In this paper, we followed Han et al. (2022) and set $M=10$ for all experiments.
+$QICE:=$ $\frac{1}{M}\sum_{m=1}^{M}|r_{m}-\frac{1}{M}|,$ (15)
+$\text{where}\ \ r_{m}=$ $\frac{1}{N}\sum_{n=1}^{N}\mathbb{I}\{\hat{y}_{n}\geq q_{\alpha/2}\}\cdot\mathbb{I}\{\hat{y}_{n}\leq q_{1-\alpha/2}\}.$
+In this paper, we followed (Han et al., 2022) and set $M=10$ for all experiments.
 3.2 Toy Examples
-To demonstrate the effectiveness of DistPred, we initially conduct experiments on $8$ toy examples as done in CARD (Han et al., 2022).
-These examples are specifically crafted with distinct statistical characteristics in their data generating functions: some have a uni-modal symmetric distribution for their error term (linear regression, quadratic regression, sinusoidal regression), others have heteroscedasticity (log-log linear regression, log-log cubic regression) or multi-modality (inverse sinusoidal regression, 8 Gaussians, full circle).
-The research demonstrates that a trained DistPred models has the capability to produce samples that closely resemble the true response variable for novel covariates.
+To demonstrate the effectiveness of DistPred, we initially conducted experiments on $8$ toy examples as done in (Han et al., 2022).
+These examples are specifically crafted with distinct statistical characteristics in their data generating functions: some have a uni-modal symmetric distribution for their error term (linear regression, quadratic regression, sinusoidal regression), while others exhibit heteroscedasticity (log-log linear regression, log-log cubic regression) or multi-modality (inverse sinusoidal regression, $8$ Gaussians, full circle).
+The research demonstrates that a trained DistPred model has the capability to produce samples that closely resemble the true response variable for novel covariates.
 Additionally, it can quantitatively match the true distribution based on certain summary statistics.
-The study visualizes scatter plots comparing real and generated data for all eight tasks in Figure 5.
+The study visualizes scatter plots comparing real and generated data for all eight tasks in Fig.
+4.
 In cases where the tasks involve unimodal conditional distributions, the interest region fills the region between the $2.5$th and $97.5$th percentiles of the generated $\hat{y}$ values.
@@
 This indicates that the advantages of DistPred mentioned earlier can be fully harnessed in distribution prediction.
-Figure 5: Scatter plot of DistPred’s regression results on $8$ toy examples.
+Figure 5: DistPred and DistPred+MCD on UCI datasets.
+Figure 6: Ablation studies of the number of samples in DistPred (a) and the number of ensembles in DistPred+MCD (b).
 3.3 UCI Regression Tasks
+Table 1: Comparison of model training and inference times (minutes) on UCI regression datasets when $K=1000$.
+Models DistPred DistPred+MCD PBP MC Dropout CARD
+Training 0.035 $\pm$ 6E-3 0.035 $\pm$ 6E-3 0.04 $\pm$ 8E-3 0.031 $\pm$ 0.01 8.14$\pm$ 0.05
+Inference 0.026 $\pm$ 4E-3 0.027 $\pm$ 5E-3 5.23 $\pm$ 0.10 4.62 $\pm$ 0.06 8.31 $\pm$ 0.17
+Figure 7: Comparison DistPred (ours) with conformal prediction methods. All results are averaged across all prediction lengths.
 For experiments conducted on real-world datasets, we utilize the same $10$ UCI regression benchmark datasets (Asuncion and Newman, 2007) and follow the experimental protocol introduced by Hernández-Lobato and Adams (2015), which has also been followed by Gal and Ghahramani (2016) and Lakshminarayanan et al.
@@
 (2022).
-The dataset information can be found in Table 5 located in Appendix B.
+The dataset information can be found in Table 7 located in Appendix A.
 […]
@@
 # Top 2 $2$ 00 $2$ $1$ $6$ $9$ $10$
-We compare DistPred with other state-of-the-art methods, including PBP (Hernández-Lobato and Adams, 2015), MC Dropout (Gal and Ghahramani, 2016), DeepEnsemble (Lakshminarayanan et al., 2017), and another deep generative model that estimates a conditional distribution sampler, GCDS (Zhou et al., 2021b), as well as diffusion model, CARD (Han et al., 2022).
+We compare DistPred with other state-of-the-art methods, including PBP (Hernández-Lobato and Adams, 2015), MC Dropout (Gal and Ghahramani, 2016), DeepEnsemble (Lakshminarayanan et al., 2017), and another deep generative model that estimates a conditional distribution sampler, GCDS (Zhou et al., 2021b), as well as the diffusion model, CARD (Han et al., 2022).
 The multiple train-test splits are applied with a 90%/10% ratio, following the same methodology as Hernández-Lobato and Adams (2015) and Han et al.
@@
 (2022), we compare the QICE of different methods on various UCI datasets.
-Additional information regarding the experimental setup for these models is available in Appendix C.
+Additional information regarding the experimental setup for these models is available in Appendix A.2.
 The experimental results, along with corresponding metrics, are presented in Table 2.
 The frequency with which each model achieves the best corresponding metric is reported in the penultimate row, while the frequency with which it achieves the top two positions is reported in the last row.
-The results demonstrate that the DistPred method outperforms existing methods, often by a considerable margin.
+The results demonstrate that DistPred outperforms existing methods, often by a considerable margin.
 It is worth noting that these impressive results are achieved in a single forward pass of the DistPred method.
-Crucially, the performance of uncertain quantization can be enhanced even further by leveraging DistPred-MCD, a hybrid approach combining DistPred and MC Dropout.
-The implementation of DistPred on UCI regression tasks follows a straightforward approach: We employ a basic multilayer perceptron (MLP) as the foundational framework, complemented by Equation 9 serving as the loss function for end-to-end training.
+Crucially, as shown in Fig.
+5, the performance of uncertain quantization can be further enhanced by leveraging DistPred+MCD, a hybrid approach that combines DistPred and MC Dropout.
+The implementation of DistPred on UCI regression tasks follows a straightforward approach: We employ a basic multilayer perceptron (MLP) as the foundational framework, complemented by Eq.
+13 serving as the loss function for end-to-end training.
 Due to the fact that DistPred necessitates solely a single forward inference, its inference speed is notably rapid.
 Table 1 presents a comparison of the training and inference speeds of mainstream models.
-It should be noted that, for a fair comparison, the implementations of various models are constructed on the same backbone and utilized the same equipment.
-It is evident that DistPred is approximately $230$ times faster in training and about $90$ times faster in inference compared to the state-of-the-art model CARD.
+It should be noted that, for a fair comparison, the implementations of various models are constructed on the same backbone and utilize the same equipment.
+It is evident that DistPred is approximately achieves at least 180x faster inference speed compared to state-of-the-art models.
 The inference speed of DistPred is slower than its training speed because it involves calculating distribution statistical metrics like QICE and PICP.
-Table 2: Comparison of model training and inference times (minutes) on UCI boston datasets.
-Models DistPred DistPred-MCD PBP MC Dropout CARD
-Training 0.035 $\pm$ 6E-3 0.035 $\pm$ 6E-3 0.04 $\pm$ 8E-3 0.031 $\pm$ 0.01 8.14$\pm$ 0.05
-Inference 0.095 $\pm$ 4E-3 0.027 $\pm$ 5E-3 5.23 $\pm$ 0.10 4.62 $\pm$ 0.06 8.31 $\pm$ 0.17
 3.4 Ablation Study of The Number of Samples and Ensembles
-We investigate the influence of the number of samples generated by DistPred, as well as the number of ensembles of DistPred-MCD, on their respective performances.
-As shown in Figure 6, with an increase in the number of output samples and ensembles, the model’s performance shows a gradual improvement, eventually reaching a point of saturation.
-Figure 6: Ablation study of the number of samples in DistPred (a) and the number of ensembles in DistPred-MCD (b).
-3.5 Time series distribution forecasting
-We extend time series forecasting (Zhou et al., 2021a; Wu et al., 2021; Zhou et al., 2022; Liu et al., 2023) from point estimation to the task of distribution prediction to infer about more statistical information about a certain moment.
-Baselines: We employ recent 10 SOTA methods for comparisons, including iTransformer Liu et al.
-(2023), PatchTST Nie et al.
-(2022), SCINet Liu et al.
-(2022a), TimesNet Wu et al.
-(2022), DLinear Zeng et al.
-(2023), FEDformer Zhou et al.
-(2022), Autoformer Wu et al.
-(2021), Informer Zhou et al.
-(2021a), LogTrans Li et al.
-(2019) and Reformer Kitaev et al.
-(2020).
-We use the same experimental setup as (Zhou et al., 2021a) and (Liu et al., 2022a) and follow the same experimental protocol as (Zhou et al., 2021a).
-Univariate results can be found in Appendix D.
-Datasets and setting: The detailed information pertaining to the datasets can be located in Appendix B.
-The models used in the experiments are evaluated over a wide range of prediction lengths to compare performance on different future horizons: 96, 192, 336 and 720.
-The experimental settings are the same for both multivariate and univariate tasks.
-We use the average of the MSE and MAE ($\frac{MSE+MAE}{2}$) to evaluate the overall performance of the model.
-It is noteworthy that DistPred provides an ensemble $\hat{Y}$ of response variables.
-Consequently, we employ the mean value of $\hat{Y}$ as the point estimate at that moment.
-Main results: The results for multivariate TS forecasting are outlined in Table 3, with the optimal results highlighted in bold and the second-best results emphasized with underlined.
-It can be found that, despite not utilizing MSE and MAE, DistPred achieves state-of-the-art performance across all datasets and prediction length configurations.
-iTransformer and PatchTST stand out as the latest models acknowledged for their exceptional average performance.
-Compared with them, the proposed DistPred demonstrates an average performance increase of 3.5% and 16.5%, respectively, achieving a substantial performance improvement.
-We provide metrics, e.g., CRPS, QICE, PICP, for comparison by the future research community.
+We investigate the influence of the number of samples generated by DistPred, as well as the number of ensembles of DistPred+MCD, on their respective performances.
+In Fig.
+6(a), we increase the number of samples of DistPred from 10 to 4000 to observe the changes in its QICE.
+In Fig.
+6(b), we increase the number of ensembles of DistPred+MCD from 1 (DistPred) to 1000 to observe the changes in its QICE.
+It can be found that with an increase in the number of output samples and ensembles, the model’s performance shows a gradual improvement.
 Table 3: Multivariate time series forecasting results on six benchmark datasets.
@@
 # Top 1 - - - 22 28 5 1 0 1 1 0 0 0 1 0 0 0 0 0 2 0
+3.5 Time Series Distribution Forecasting
+We extend time series forecasting (Zhou et al., 2021a; Wu et al., 2021; Zhou et al., 2022; Liu et al., 2023) from point estimation to the task of distribution prediction to infer about more statistical information about a certain moment.
+Baselines: We employ recent 10 SOTA methods for comparisons, including iTransformer Liu et al.
+(2023), PatchTST Nie et al.
+(2022), SCINet Liu et al.
+(2022a), TimesNet Wu et al.
+(2022), DLinear Zeng et al.
+(2023), FEDformer Zhou et al.
+(2022), Autoformer Wu et al.
+(2021), Informer Zhou et al.
+(2021a), LogTrans Li et al.
+(2019) and Reformer Kitaev et al.
+(2020).
+DistPred employs the same network architecture as iTransformer.
+We use the same experimental setup as (Zhou et al., 2021a) and (Liu et al., 2022a) and follow the same experimental protocol as (Zhou et al., 2021a).
+Univariate results can be found in Appendix E.
+Datasets and Setting: The detailed information pertaining to the datasets can be located in Appendix A.
+The models Liang et al.
+(2024) used in the experiments are evaluated over a wide range of prediction lengths to compare performance on different future horizons: 96, 192, 336, and 720.
+The experimental settings are the same for both multivariate and univariate tasks.
+We use the average of the MSE and MAE ($\frac{MSE+MAE}{2}$) to evaluate the overall performance of the model.
+It is noteworthy that DistPred provides an ensemble $\hat{Y}$ of response variable.
+Consequently, we employ the mean value of $\hat{Y}$ as the point estimate at that moment.
+Main Results: The results for multivariate TS forecasting are outlined in Table 3, with the optimal results highlighted in bold and the second-best results emphasized with underlined.
+It can be found that, despite not utilizing MSE and MAE, DistPred achieves state-of-the-art performance across all datasets and prediction length configurations.
+iTransformer and PatchTST stand out as the latest models acknowledged for their exceptional average performance.
+Compared with them, the proposed DistPred demonstrates an average performance increase of 3.5% and 16.5%, respectively, achieving a substantial performance improvement.
+We provide metrics, e.g., CRPS, QICE, PICP, for comparison by the future research community.
+Note that for long time series, computing these metrics on the entire testset can be very time-consuming and may lead to out-of-memory issues.
+Therefore, we propose calculating these metrics for each batch and then averaging the results.
+3.6 Comparison with Conformal Prediction
+Since conformal prediction offers an alternative approach for distribution prediction, we compare DistPred with conformal prediction methods on the time series forecasting tasks, including CQR (Romano et al., 2019), EnbPI (Xu and Xie, 2021), SPCI (Xu and Xie, 2023).
+Since PICP is the primary metric used by these methods, we adhere to their convention in this work.
+As shown in Fig.
+7, DistPred achieves better performance than conformal prediction methods on all datasets and prediction lengths, with an average PICP improvement of 4.5%.
+This further underscores the high competitiveness of DistPred.
+Table 4: Stability experiments of loss function (Eq. 13).
+1% 25% 50% 75% 100%
+Train Loss 0.21$\pm$0.04 0.174$\pm$0.03 0.165$\pm$0.01 0.162$\pm$0.01 0.16$\pm$0.007
+Valid Loss 0.17$\pm$0.05 0.155$\pm$0.04 0.151$\pm$0.02 0.15$\pm$0.02 0.15$\pm$0.01
+Table 5: KL-divergence, skewness, and kurtosis of the predictive response variable to the standard Gaussian on the ETTm2 datasets under input-96-predict-96 with $K=100$.
+Variates 1 2 3 4 5 6 7
+KL-Div 0.11 0.1 inf 0.13 0.03 inf 0.05
+Skewness -0.42 0.41 -0.47 0.16 -0.03 -0.26 -0.72
+Kurtosis 0.77 0.09 1.93 0.31 0.85 1.95 1.77
+3.7 Stability of The Loss Function
+We included detailed discussions and experiments on the convergence and stability of the proposed method.
+As shown in Table 4, the experiments show that the training and valid loss converge smoothly across multiple runs, with no evidence of instability.
+Figure 8: QQ-plots and histograms of response variable.
 3.8 Visualization of The Predictive Distribution
-The predictive distribution of DistPred is visualized in Figure 7.
-It can be observed that DistPred gives an ensemble of predictions (only top 10 are presented in the left subplot).
+We conducted thorough statistical analyses, including QQ-plots, histograms, and metrics like KL-divergence, skewness, and kurtosis, to validate the empirical distribution of the predictive response variable.
+KL-divergence, skewness, and kurtosis of the predictive response variable are shown in Table 5, and QQ-plots and histograms are shown in Fig.
+8.
+These statistical analyses shows that the empirical distribution exhibits notable characteristics, such as skewness and heavy tails, etc., which are typical of many real-world datasets.
+Furthermore, the predictive distribution of DistPred is visualized in Fig.
+9.
+It can be observed that DistPred provides an ensemble of predictions (only the top 10 are presented in the left subplot).
 Given all predictive ensemble values, the model can estimate the distribution of the response variable.
-Therefore, we can calculate confidence intervals at different levels, as shwon in the right subplot of Figure 7.
-Figure 7: Visualization of the prediction results and the confidence intervals.
-4 Conclusion
-In this paper, we propose a novel method called DistPred, which is a distribution-free probabilistic inference method for regression and forecasting tasks.
-We transform proper scoring rules that measure the discrepancy between the predicted distribution and the target distribution into a differentiable discrete form and use it as a loss function to train the model end-to-end.
-This allows the model to sample numerous samples in a single forward pass to estimate the potential distribution of the response variable.
-We also propose a hybrid method called DistPred-MCD, which combines DistPred with MC Dropout to further improve the performance of uncertain quantization.
-Experimental results demonstrate that DistPred outperforms existing methods, often by a considerable margin.
-We also extend time series forecasting from point estimation to distribution prediction and achieve state-of-the-art performance on multivariate and univariate time series forecasting tasks.
-In the future, we plan to extend DistPred to other tasks, such as classification and reinforcement learning.
-Appendix A Related work
+Consequently, we can calculate confidence intervals at different levels, as shown in the right subplot of Fig.
+9.
+Figure 9: Visualization of the prediction results and the confidence intervals with setting input-96-predict-96 on the Ettm2 dataset.
+(a) The left subplot shows the ensemble of predictions with $K=100$.
+By utilizing subgraph (a), we can directly obtain the confidence intervals for subgraph (b), e.g, confidence intervals at 99%, 99.5% and 99.9% levels.
+4 Related Work
 In supervised learning contexts, the endeavor to characterize the conditional distribution $p(y|x)$ beyond merely the conditional mean $\mathop{\mathbb{E}}[y|x]$ via deep neural networks has been a focal point of existing research efforts.
-These endeavors primarily concentrate on quantifying predictive uncertainty, with several approaches having been proposed.
-Bayesian neural networks (BNNs) represent one such approach, aiming to capture such uncertainty by positing distributions over network parameters, thereby encapsulating the model’s plausibility given the available data (Blundell et al., 2015; Hernández-Lobato and Adams, 2015; Gal and Ghahramani, 2016; Kingma et al., 2015; Tomczak et al., 2021).
-Another avenue is represented by Kendall and Gal (2017), which not only addresses uncertainties in model parameters but also incorporates an additive noise term into the neural network output to encompass uncertainties in model outputs.
+These endeavors primarily concentrate on quantifying uncertainty, with several approaches having been proposed.
+In regression and forecasting tasks, predicting the underlying distribution of the response variable is pivotal.
+Traditional approaches (Bishop, 1994; Greene, 2003; Salinas et al., 2020; He et al., 2020) often rely on the assumption that the response variable conforms to a prior distribution, with the goal of estimating specific statistics derived from this distribution.
+Commonly, these methods transform the challenge of distribution prediction and uncertainty quantification into the prediction of statistical parameters such as the mean and variance, under the presumption of a known continuous distribution.
+For example, MDNs (Bishop, 1994) impose a predefined distribution—typically Gaussian—weighted by certain parameters to approximate the final distribution.
+Similarly, heteroscedastic regression (Greene, 2003) predicts uncertainty by modeling the variability of residuals as a function of independent variables.
+DeepAR (Salinas et al., 2020), another notable approach, assumes a Gaussian distribution for the response variable, thereby leveraging the GaussianNLLLoss (Nix and Weigend, 1994) to directly optimize its mean and variance.
+While these methods offer computational efficiency by simplifying predictions to statistical parameters, their reliance on strong distributional assumptions often limits their ability to capture the true underlying distribution, potentially leading to suboptimal performance.
+Conformal prediction presents an alternative framework for distribution prediction, diverging from traditional parametric approaches.
+In their study, the authors in (Vovk et al., 2017, 2018) introduced a random prediction system and proposed a nonparametric prediction method grounded in conformal assumptions.
+By integrating conformal prediction with quantile regression in (Romano et al., 2019; Xu and Xie, 2021, 2023), they developed a method for constructing prediction intervals for the response variable.
+However, the practical application of conformal prediction is not without limitations.
+Its effectiveness is often constrained by the assumption of exchangeability of residuals, which may not hold in all contexts, particularly in the presence of temporal dependencies.
+This limitation can lead to less reliable prediction intervals when applied to non-independent and identically distributed (non-i.i.d.) data, thereby challenging its robustness in real-world scenarios where data often exhibit complex dependencies.
+Uncertainty quantification is a method that indirectly reflects the potential distribution of the response variable.
+BNNs represent one such approach, aiming to capture such uncertainty by positing distributions over network parameters, thereby encapsulating the model’s plausibility given the available data (Blundell et al., 2015; Hernández-Lobato and Adams, 2015; Gal and Ghahramani, 2016; Kingma et al., 2015; Tomczak et al., 2021).
+Another avenue is represented by Kendall and Gal (2017), which not only addresses uncertainties in model parameters but also incorporates an additive noise term into the outputs to encompass uncertainties.
 In parallel, ensemble-based methodologies (Lakshminarayanan et al., 2017; Liu et al., 2022b) have emerged to address predictive uncertainty.
-These methods involve amalgamating multiple neural networks with stochastic outputs, thereby providing a comprehensive approach to uncertainty quantification.
-Furthermore, the neural processes’ family (Garnelo et al., 2018b, a; Kim et al., 2019; Gordon et al., 2020) has introduced a suite of models tailored to capturing predictive uncertainty in a manner that extends beyond the distribution of available data, particularly tailored for few-shot learning scenarios.
+These methods involve amalgamating multiple neural networks with stochastic outputs.
+Furthermore, the neural processes’ family (Garnelo et al., 2018b, a; Kim et al., 2019; Gordon et al., 2020) has introduced a suite of models tailored to capturing predictive uncertainty in a manner that extends beyond the distribution of available data, particularly tailored for few-shot learning.
 The aforementioned models have predominantly operated under the assumption of a parametric form in $p(y|x)$, typically adopting a Gaussian distribution or a mixture of Gaussians.
@@
 Nevertheless, it is imperative to note that these models entail protracted training processes and computationally demanding inference procedures.
-Appendix B Dataset
-B.1 Commonly used TS datasets
+5 Conclusion
+In this paper, we propose a novel method named DistPred, which is a distribution-free probabilistic inference approach for regression and forecasting tasks.
+We transform proper scoring rules that measure the discrepancy between the predicted distribution and the target distribution into a differentiable discrete form and use it as a loss function to train the model end-to-end.
+This allows the model to sample numerous samples in a single forward pass to estimate the potential distribution of the response variable.
+Experimental results demonstrate that DistPred outperforms existing methods, often by a considerable margin.
+We also extend time series forecasting from point estimation to distribution prediction and achieve state-of-the-art performance on multivariate and univariate time series forecasting tasks.
+In the future, we plan to extend DistPred to other tasks, such as classification and reinforcement learning.
+Appendix A Dataset and Implementation
+A.1 Commonly Used TS Datasets
+The information of the experiment datasets used in this paper are summarized as follows: (1) Electricity Transformer Temperature (ETT) dataset Zhou et al.
+(2021a), which contains the data collected from two electricity transformers in two separated counties in China, including the load and the oil temperature recorded every 15 minutes (ETTm) or 1 hour (ETTh) between July 2016 and July 2018.
+(2) Electricity (ECL) dataset  [^1] https://archive.ics.uci.edu/ml/datasets/ElectricityLoadDiagrams20112014 collects the hourly electricity consumption of 321 clients (each column) from 2012 to 2014.
+(3) Exchange Lai et al.
+(2018) records the current exchange of 8 different countries from 1990 to 2016.
+(4) Traffic dataset  [^2] http://pems.dot.ca.gov records the occupation rate of freeway system across State of California measured by 861 sensors.
+(5) Weather dataset  [^3] https://www.bgc-jena.mpg.de/wetter records every 10 minutes for 21 meteorological indicators in Germany throughout 2020.
+(6) Solar-Energy Lai et al.
+(2018) documents the solar power generation of 137 photovoltaic (PV) facilities in the year 2006, with data collected at 10-minute intervals.
+(7) The PEMS dataset Liu et al.
+(2022a) comprises publicly available traffic network data from California, collected within 5-minute intervals and encompassing 358 attributes.
+(8) Illness (ILI) dataset  [^4] https://gis.cdc.gov/grasp/fluview/fluportaldashboard.html describes the influenza-like illness patients in the United States between 2002 and 2021, which records the ratio of patients seen with illness and the total number of the patients.
+The detailed statistics information of the datasets is shown in Table 6, and the dataset information in terms of their size and number of features is summarized in Table 7.
+A.2 Implementation Details
+The model undergoes training utilizing the ADAM optimizer Kingma and Ba (2015) and minimizing the Mean Squared Error (MSE) loss function.
+The training process is halted prematurely, typically within 10 epochs.
+The DistPred architecture solely comprises the embedding layer and backbone architecture, devoid of any additional introduced hyperparameters.
+During model validation, two evaluation metrics are employed: CRPS, QICE, PICP, MSE and MAE.
+Given the potential competitive relationship between the two indicators, MSE and MAE, we use the average of the two ($\frac{MSE+MAE}{2}$) to evaluate the overall performance of the model.
 Table 6: Details of the seven TS datasets.
@@
 Illness 966 7 7d
-The information of the experiment datasets used in this paper are summarized as follows: (1) Electricity Transformer Temperature (ETT) dataset Zhou et al.
-(2021a), which contains the data collected from two electricity transformers in two separated counties in China, including the load and the oil temperature recorded every 15 minutes (ETTm) or 1 hour (ETTh) between July 2016 and July 2018.
-(2) Electricity (ECL) dataset  [^1] https://archive.ics.uci.edu/ml/datasets/ElectricityLoadDiagrams20112014 collects the hourly electricity consumption of 321 clients (each column) from 2012 to 2014.
-(3) Exchange Lai et al.
-(2018) records the current exchange of 8 different countries from 1990 to 2016.
-(4) Traffic dataset  [^2] http://pems.dot.ca.gov records the occupation rate of freeway system across State of California measured by 861 sensors.
-(5) Weather dataset  [^3] https://www.bgc-jena.mpg.de/wetter records every 10 minutes for 21 meteorological indicators in Germany throughout 2020.
-(6) Solar-Energy Lai et al.
-(2018) documents the solar power generation of 137 photovoltaic (PV) facilities in the year 2006, with data collected at 10-minute intervals.
-(7) The PEMS dataset Liu et al.
-(2022a) comprises publicly available traffic network data from California, collected within 5-minute intervals and encompassing 358 attributes.
-(8) Illness (ILI) dataset  [^4] https://gis.cdc.gov/grasp/fluview/fluportaldashboard.html describes the influenza-like illness patients in the United States between 2002 and 2021, which records the ratio of patients seen with illness and the total number of the patients.
-The detailed statistics information of the datasets is shown in Table 4.
-The dataset information in terms of their size and number of features is summarized in Table 5.
 Table 7: Dataset size ($N$ observations, $P$ features) of UCI regression tasks.
@@
 $(N,P)$ $(506,13)$ $(1030,8)$ $(768,8)$ $(8192,8)$ $(11,934,16)$ $(9568,4)$ $(45,730,9)$ $(1599,11)$ $(308,6)$ $(515,345,90)$
-Table 6: Batch size settings of UCI regression tasks across different models.
-PBP MC Dropout Deep Ensembles GCDS CARD (ours)
-Boston $32$ $32$ $32$ $32$ $32$
-Concrete $32$ $32$ $32$ $32$ $32$
-Energy $32$ $32$ $32$ $32$ $32$
-Kin8nm $64$ $32$ $64$ $64$ $64$
-Naval $64$ $32$ $64$ $64$ $64$
-Power $64$ $64$ $64$ $64$ $64$
-Protein $100$ $256$ $100$ $256$ $256$
-Wine $32$ $32$ $32$ $32$ $32$
-Yacht $32$ $32$ $32$ $32$ $32$
-Year $256$ $256$ $100$ $256$ $256$
-B.2 Distribution free vs. distribution related
+Appendix B Comparisons of Model Efficiency in Time Series Forecasting
+We evaluate the inference time and size of DistPred (mean only) in comparison to other Transformer-based models, as shown in Fig.
+10.
+It can be observed that DistPred’s runtime, model size, and memory usage do not significantly increase compared to other models.
+Figure 10: QQ-plots and histograms of response variable.
+Appendix C Distribution Free v.s. Distribution Related
 If we assume that the response variable follows a continuous distribution, as done in (Salinas et al., 2020) where $\hat{y}$ is assumed to be followed Gaussian distribution, we can provide an analytical formula for the Gaussian likelihood.
@@
 $\mathcal{L}_{G}(\hat{y}|\mu,\sigma)=(2\pi\sigma^{2})^{-\frac{1}{2}}exp(-(\hat{y}-\mu)^{2}/(2\sigma^{2})).$ (16)
-Then, we can train DistPred utilizing Equation 13 as the loss function. Currently, DistPred exclusively provides the mean and variance of the response variable.
+Then, we can train DistPred utilizing Eq. 16 as the loss function. As shown in Table 8, DistPred exclusively provides the mean and variance of the response variable.
 Table 8: Comparison of DistPred trained with Gaussia distribution.
@@
 DistPred-G 4391.80 7.61 1.73 2329.05 276.51 6.35 177.92 16.52 1.72 229.58 2.27 0.17
-Appendix C Implementation details
-The model undergoes training utilizing the ADAM optimizer Kingma and Ba (2015) and minimizing the Mean Squared Error (MSE) loss function.
-The training process is halted prematurely, typically within 10 epochs.
-The DistPred architecture solely comprises the embedding layer and backbone architecture, devoid of any additional introduced hyperparameters.
-During model validation, two evaluation metrics are employed: CRPS, QICE, PICP, MSE and MAE.
-Given the potential competitive relationship between the two indicators, MSE and MAE, we use the average of the two ($\frac{MSE+MAE}{2}$) to evaluate the overall performance of the model.
-Appendix D Univariate time series forecasting
+Appendix D DistPred for Time Series with Missing Values
+The proposed DistPred is equipped to handle time series with missing values scenarios, as both imputation and forecasting are done in a similar manner.
+Our experiments have demonstrated that DistPred remains SOTA even when a substantial portion of the data is missing.
+Specifically, in cases where 80% of the PhysioNet2012 data points were absent, the model still maintained competitive performance, as shown in Table 9.
+Table 9: Comparison of DistPred with other SOTA models when time series with missing values.
+Model M-RNN GP-VAE BRITS USGAN CSDI TimesNet Transformer SAITS DistPred
+MSE 0.864$\pm$0.002 0.433$\pm$0.011 0.325$\pm$0.002 0.306$\pm$0.001 0.260$\pm$0.05 0.272$\pm$0.006 0.225$\pm$0.002 0.218$\pm$0.002 0.204$\pm$0.002
+MAE 0.674$\pm$0.001 0.4$\pm$0.007 0.246$\pm$0.001 0.25$\pm$0.001 0.211$\pm$0.003 0.266$\pm$0.007 0.209$\pm$0.002 0.202$\pm$0.002 0.196$\pm$0.002
+Appendix E Univariate Time Series Forecasting
 The full results for univariate TS forecasting are presented in Table 10.
@@
 $1^{\text{st}}$ Count - - - 34 37 3 1 0 0 1 1 2 1 0 0
-Appendix E Limitations of DistPred
-DistPred is not without its limitations.
-Firstly, the model is not designed to handle categorical data, and as such, it is not suitable for classification tasks.
-Secondly, the model is not designed to handle time series data with missing values.
-Thirdly, the model is not designed to handle multivariate time series data.
-Finally, the model is not designed to handle time series data with irregular time intervals.